### v1

# Communication-Efficient Distributed Training for Collaborative Flat Optima Recovery in Deep Learning

Abstract
We study centralized distributed data parallel training of deep neural networks (DNNs), aiming to improve the trade-off between communication efficiency and model performance of the local gradient methods. To this end, we revisit the flat-minima hypothesis, which suggests that models with better generalization tend to lie in flatter regions of the loss landscape. We introduce a simple, yet effective, sharpness measure, Inverse Mean Valley, and demonstrate its strong correlation with the generalization gap of DNNs. We incorporate an efficient relaxation of this measure into the distributed training objective as a lightweight regularizer that encourages workers to collaboratively seek wide minima. The regularizer exerts a pushing force that counteracts the consensus step pulling the workers together, giving rise to the Distributed Pull-Push Force (DPPF) algorithm. Empirically, we show that DPPF outperforms other communication-efficient approaches and achieves better generalization performance than local gradient methods and synchronous gradient averaging, while significantly reducing communication overhead. In addition, our loss landscape visualizations confirm the ability of DPPF to locate flatter minima. On the theoretical side, we show that DPPF guides workers to span flat valleys, with the final valley width governed by the interplay between push and pull strengths, and that its pull-push dynamics is self-stabilizing. We further provide generalization guarantees linked to the valley width and prove convergence in the non-convex setting.
1 Introduction
Deep Learning (DL) nowadays is applied to a plethora of complex learning problems across vision dosovitskiy2020image ; zou2023object ; bojarski2016end , language vaswani2017attention ; kenton2019bert ; radford2019language , and multi-modal domains mao2017deepart . DNNs are trained using the Stochastic Gradient Descent (SGD) algorithm bottou-98x or its variants amari1993backpropagation ; ruder2016overview ; kingma2014adam ; duchi2011adaptive . The size of these networks often scales with both the training dataset DBLP:conf/kdd/ChenWTWC16 ; imagenet ; bahri2024explaining ; Zhai_2022_CVPR and task complexity DBLP:journals/corr/abs-1810-04805 . Modern state-of-the-art models contain billions of parameters Zhai_2022_CVPR ; MegatronLM ; mann2020language ; achiam2023gpt and are trained on massive datasets using thousands of GPUs over weeks Auto ; Amazon , incurring enormous computational costs Money . To speed up training at this scale, data parallelism is widely adopted, where data is distributed across multiple devices stich2019local ; pytorch_distributed ; parallel_and_distributed . In this work, we consider data-parallel training with a centralized server.
In data parallelism, a straightforward and widely used approach to scaling DNN training across multiple devices is to split the dataset among workers, where a worker refers to a single hardware unit (CPU, GPU, TPU, etc.). Each worker computes gradients on its data shard, and these gradients are then aggregated (for example, by averaging) and applied synchronously to update model replicas across all devices mcdonald2009efficient ; pytorch_distributed . While this gradient-sharing scheme is effective, it introduces a major communication bottleneckharlap2018pipedream , as it requires synchronization among all workers at every iteration.
An alternative distributed training strategy enables communication efficiency by allowing workers to take several local gradient steps independently before periodically aligning with a consensus variable—either through hard resets stich2019local or soft pulling zhang2015deep . Although this reduces the number of communication rounds compared to gradient averaging, it often results in inferior performance compared to fully synchronous training yu2019parallel ; ortiz2021trade .
Our goal in this work is to develop a mechanism that enables communication-efficient training methods to match/exceed the performance of synchronous gradient averaging, without incurring significant computational overhead. To this end, we revisit the flatness hypothesis in the literature, which suggests that models with better generalization tend to converge to flatter regions of the loss landscape hochreiter1997flat ; keskar2016large ; bisla2022low ; SAM . Our key contributions can be listed as follows:
  • We introduce a new sharpness measure, Inverse Mean Valley (Inv. MV), which shows strong correlation with the generalization gap in DNNs trained using communication-efficient methods and outperforms existing sharpness metrics in comparative evaluations.
  • We propose a lightweight relaxation of the Inv. MV measure that can be efficiently integrated into the training objective, inducing a push force that counteracts the periodic consensus steps in communication-efficient methods. This gives rise to our algorithm, Distributed Pull-Push Force (DPPF), where the push component keeps workers apart, preventing collapse, avoiding convergence to narrow valleys, and positioning them near the boundaries of wide, flat regions in the loss landscape.
  • We theoretically characterize the interplay between the pull and push forces in DPPF and show how they govern the final distance between the workers and their average. We also provide generalization guarantees tied to this distance, along with a non-convex convergence analysis. Finally, we provide theoretical insights into the workers’ relative positioning at convergence.
  • We experimentally validate our approach on standard benchmark datasets and architectures, conduct extensive ablation studies to analyze its underlying mechanisms, and confirm its ability to recover wide minima through loss landscape visualizations.
2 Related Work
Input : Pull force $\alpha\in(0,1]$, comm. period $\tau$, learning rate $\eta$
Initialize $x_{1},...,x_{M}$, $t_{1}=\ldots=t_{M}=0$, and data shards $\Psi_{1},...,\Psi_{M}$
for each worker $m$ in parallel do
  while not converged do
     Draw batch $\xi_{m}\in\Psi_{m}$
     $x_{m}\leftarrow x_{m}-\eta\nabla f(x_{m};\xi_{m})$
     $t_{m}\leftarrow t_{m}+1$
     if $t_{m}\bmod\tau=0$ then
        Obtain $x_{C}$
        $x_{m}\leftarrow(1-\alpha)x_{m}+\alpha x_{C}$
        if push then
           $x_{m}\leftarrow x_{m}+\lambda\frac{x_{m}-x_{A}}{\|x_{m}-x_{A}\|}$
           $x_{A}=\frac{1}{M}\sum_{i=1}^{M}x_{i}$
        end if
     end if
  end while
end for
return $x_{A}$
Algorithm 1 Data Parallel Training
Data Parallel Training of DNNs A widely adopted approach in data-parallel training requires that each worker computes a stochastic gradient on its local data, aggregates these gradients via averaging, and updates the model using the aggregated gradient. This scheme—variously called Parallel Mini-Batch dist_mini_batch , All-Reduce lin2018don , or DistributedDataParallel (DDP) pytorch_distributed —is communication-intensive because gradients must be exchanged every iteration harlap2018pipedream . To reduce communication costs, local-update schemes let each worker take several SGD steps before a global average. LocalSGD stich2019local ; zhang2016parallel ; zhou2018convergence ; lian2015asynchronous exemplifies this idea and is also employed in large-scale DNN training povey2014parallel ; su2015experiments ; chen2016scalable . Although it enables communication efficiency, performance degrades when either the communication interval or worker count grows stich2019local ; yu2019parallel ; ortiz2021trade ; adding a global momentum term partially offsets this loss wangslowmo . Another way to improve model performance while training with longer communication periods was explored in the past work. Theory shows that the gradient noise from local updates can boost generalization: LocalSGD outperforms DDP when the learning rate is small and training is long whyandwhenlocalsgd . Extending this analysis, the quadratic synchronization rule (QSR) adaptively updates the communication period inversely proportional to the learning rate, yielding improvement in model generalization while being communication-efficient guquadratic . Additionally, it has been shown that gossip communication in the decentralized setting can also improve test accuracy, contradicting the notion that decentralization hurts generalization zhu2023decentralized .
(a) ResNet-18, CIFAR-10, 4W
(b) PyramidNet, CIFAR-100, 8W
(c) ResNet-50, ImageNet, 4W
Figure 1: Communication volume (ratio between the number of communication rounds till convergence to the total number of local iterations; lower is better) vs. test error (%) for baseline distributed training methods (LocalSGD, QSR and DDP SGD) and our approach, DPPF. (W: Worker)
Since full consensus can hinder local exploration, several soft-consensus methods have been proposed to balance synchronization and exploration. Elastic Averaging SGD (EASGD) softly pulls workers to a moving-average center to preserve exploration zhang2015deep . Leader SGD (LSGD) pulls all workers toward the one with the lowest loss to accelerate convergence teng2019leader , while GRAWA takes a flatness-aware approach by weighting workers inversely with their gradient norms, encouraging optimization toward flatter regions of the loss landscape dimlioglu2024grawa .
Flatness-Aware Optimizers The concept of flat minima dates back to hochreiter1994simplifying , and has since inspired extensive research on the geometry of DNN loss landscapes goodfellow2014qualitatively ; dauphin2014identifying ; baldassi2015subdominant ; li2018visualizing and its link to generalization keskar2016large ; jastrzkebski2017three ; andriushchenko2023modern . While many studies associate sharpness with generalization, relatively few optimization methods are explicitly designed to seek flat minima. Some examples include Entropy-SGD chaudhari2019entropy , which smooths the loss surface via local entropy, and Low-Pass Filtering SGD bisla2022low , which optimizes a Gaussian-smoothed objective function. A widely popularized method is Sharpness-Aware Minimization (SAM) SAM ; kwon2021asam ; zhuang2022surrogate ; li2024enhancing , which seeks parameters that lie in neighborhoods having uniformly low loss via a min-max optimization. In contrast to these techniques, Random Weight Perturbation (RWP) li2022efficient ; lirevisiting applies random directional perturbations. Another notable approach is Stochastic Weight Averaging (SWA) izmailovaveraging , where averaging SGD iterates along with using a cyclical or constant learning rate yields flatter solutions and improved generalization.
Note: Review of additional related works can be found in Section A of the Appendix.
Motivating Plot Figure 1 highlights the problem with existing baselines that struggle to sustain good performance and remain communication efficient at the same time. This challenge motivates our work. Our proposed method, allows a better tradeoff between communication and performance.
3 Preliminary
We consider a distributed data-parallel setup for training a deep neural network, where each worker $m$ has access to a disjoint data shard $\mathcal{D}_{m}$, with samples assumed to be independent and identically distributed. The goal is to collaboratively optimize a shared model vector $\mathbf{x}$ that minimizes the global training loss, formally defined in Equation 1. Here, $M$ is the number of workers, $f$ is a non-convex objective, and $F_{m}(x;\xi)$ denotes the stochastic loss observed by worker $m$. Each worker receives independent, unbiased stochastic gradients of its local objective. The generic data-parallel training procedure for optimizing this objective is outlined in Algorithm 1.
$\min_{x\in\mathbb{R}^{d}}\frac{1}{M}\sum_{m=1}^{M}f_{m}(\mathbf{x})\quad\text{where}\quad f_{m}(x)=\mathbb{E}_{\xi\sim\mathcal{D}_{m}}[F_{m}(x;\xi)]$ (1)
In particular, Algorithm 1 begins by randomly initializing the model and distributing it to all workers. The dataset is partitioned into non-overlapping shards, and each worker trains its local model using an optimizer such as SGD. Every $\tau$ iterations, workers synchronize and compute a consensus variable $x_{C}$, which can take various forms: the average of worker parameters $x_{A}$, a time-space moving average as in EASGD zhang2015deep , the leader’s parameters as in LSGD teng2019leader , or a weighted average as in MGRAWA dimlioglu2024grawa . The consensus update is then performed by interpolating between each worker’s parameters $x_{m}$ and the consensus variable $x_{C}$. When $x_{C}=x_{A}$, $\tau=1$, and $\alpha=1$, the procedure reduces to standard gradient averaging (i.e., DDP), assuming buffer synchronization. When $x_{C}=x_{A}$, $\tau>1$, and $\alpha=1$, it becomes equivalent to LocalSGD. When $\alpha<1$, workers are not forced to exactly match $x_{C}$ but are instead softly pulled toward it—a process we refer to as soft consensus. Upon completion, the algorithm returns the averaged worker parameters. The pushing mechanism, captured also in the pseudo-code, which is characteristic of DPPF, will be explained later.
4 Mean Valley Measure
In this section, we introduce the Mean Valley (MV) measure, a simple yet effective metric for quantifying the flatness of minima after full convergence in distributed training methods with independent local gradient steps. At a high level, MV estimates the average diameter of the valley surrounding the converged worker parameters. The procedure for computing MV begins with obtaining the average model $x_{A}$ and evaluating the training loss at that point $f(x_{A},\mathcal{D}_{train})$. For each worker $m$, the unit vector $\delta_{m}$ pointing from $x_{A}$ to $x_{m}$ is then calculated. A line search is performed along this direction to find the point $x_{m}^{b}$ where the average loss increases by a factor of $\kappa$ ($\kappa>1$), identifying $x_{m}^{b}$ as the valley boundary in that direction. Finally, the MV measure outputs the average distance between each boundary point $x_{m}^{b}$ and the center $x_{A}$. Mathematically, this procedure becomes:
$x_{A}=\frac{1}{M}\sum_{m=1}^{M}x_{m},\hskip 5.69054pt$ $\mathcal{L}_{A}=f\left(x_{A},\mathcal{D}_{train}\right),\hskip 5.69054pt\delta_{m}=\frac{x_{m}-x_{A}}{||x_{m}-x_{A}||_{2}}$
$x_{m}^{b}=x_{A}+\beta_{m}\delta_{m}\hskip 5.69054pt\text{s.t.}\hskip 5.69054ptf$ $\left(x_{m}^{b},\mathcal{D}_{train}\right)\approx\kappa\mathcal{L}_{A}\Rightarrow\text{MV}=\frac{1}{M}\sum_{m=1}^{M}||x_{m}^{b}-x_{A}||_{2}.$
Notice that $\beta_{m}$ is the distance we must move from $x_{A}$ along $\delta_{m}$ to reach the $\kappa$-loss contour, in particular, $\beta_{m}=\|x_{m}^{b}-x_{A}\|_{2}$. Thus, MV measure is the mean boundary radius of the valley around $x_{A}$, along the worker directions. For a fixed $\kappa>1$, a larger MV indicates lower average directional curvature (i.e., a wider minimum).
4.1 Theoretical Justification of the Mean Valley Measure
For mathematical brevity, we express $f(x;\mathcal{D}_{train})$ simply as $f(x)$. We start by writing the second-order Taylor approximation of the $f(x_{m}^{b})$ centered around $x_{A}$ (below $H$ denotes the Hessian):
$f(x_{m}^{b})=f(x_{A}+\beta_{m}\delta_{m})\approx f(x_{A})+\langle\beta_{m}\delta_{m},\nabla f(x_{A})\rangle+\frac{\beta_{m}^{2}}{2}\delta_{m}^{T}H(x_{A})\delta_{m}$
Since the analysis is carried out at the convergence, we assume full batch gradient to be close to 0, i.e. $\nabla f(x_{A})\approx\mathbf{0}$. This omits the middle term in the approximation, hence, we are left with:
$f(x_{m}^{b})=f(x_{A}+\beta_{m}\delta_{m})\approx f(x_{A})+\frac{\beta_{m}^{2}}{2}\delta_{m}^{T}H(x_{A})\delta_{m}.$
By construction, $f(x_{m}^{b})\approx\kappa f(x_{A})$, thus we can write:
$(\kappa-1)f(x_{A})\approx\frac{\beta_{m}^{2}}{2}\delta_{m}^{T}H(x_{A})\delta_{m}\quad\rightarrow\quad\beta_{m}\approx\sqrt{\frac{2(\kappa-1)f(x_{A})}{\delta_{m}^{T}H(x_{A})\delta_{m}}}$
For a reported $f(x_{A})$ and a fixed $\kappa$ value, the final expression relates $\beta_{m}$ to the directional curvature along $\delta_{m}$; specifically, $\beta_{m}\propto(\delta_{m}^{T}H(x_{A})\delta_{m})^{-0.5}$. Thus, large MV values arise from flatter directions. Since MV outputs $\frac{1}{M}\sum_{m=1}^{M}\beta_{m}$, it effectively captures the average Hessian-based curvature along the directions dictated by the workers. Moreover, with more diverse workers, MV can sample more curvature information.
4.2 Comparison with Other Sharpness Measures
Although MV is inherently a flatness measure, for consistent comparison with existing sharpness metrics, we define the Inverse Mean Valley (Inv. MV) by taking the additive inverse of MV so that the larger values indicate sharper minima. We compare Inv. MV with seven sharpness measures drawn from the literature. To quantify the correlation between sharpness measures and the generalization gap (validation--train error (%)), we use the Kendall rank correlation coefficient. We conduct this analysis by training ResNet-style resnet models on the CIFAR-10 cifar10 dataset under two settings with and without augmentation: (1) single-worker training (i.e., without distributed training), and (2) distributed training using the parameter-sharing method EASGD with 444 workers. To generate minima with different geometries, we vary several hyperparameters, model capacity, and repeat each configuration across three random seeds (see Section B.1 in the Appendix for more details on the experiments, measures, and sensitivity analysis of Inv. MV on $\kappa$).
Table 1: Kendall rank coefficients calculated between the generalization gap and the sharpness measures.
Single Worker EASGD
Measures w/ Aug. w/o Aug. w/ Aug. w/o Aug.
Shannon Ent. 0.695 0.575 -0.213 -0.161
$\epsilon$-sharpness 0.784 0.799 0.254 0.472
Fisher-Rao 0.799 0.735 0.665 0.101
LPF 0.730 0.738 0.074 0.553
$\lambda_{\max}(H)$ 0.773 0.799 0.444 0.166
[…]
$||H||_{\text{frob}}$ 0.796 0.787 0.510 0.170
Inv. MV ($\kappa=2$) NA NA 0.616 0.485
We present the correlation results in Table 1. As can be seen, the sharpness measures that have the strongest correlation with the generalization in the single-worker setting are no longer the best performers in the distributed setting. Similarly, a notable finding is that the strength of the correlations varies between scenarios with and without augmentations, e.g., the Fisher-Rao metric is well-correlated with the generalization gap for the scenario with augmentations, but without augmentations, it is one of the worst metrics. The table finally demonstrates a strong correlation of Inv. MV with the generalization gap (second-best metric), as well as it exhibits consistent behavior in settings with and without augmentations as opposed to the rest of the measures.
5 Distributed Pull-Push Force Algorithm
Motivated by the strong correlation between the Inv. MV measure and the generalization gap in DNNs trained with local-gradient methods, we propose incorporating it as a regularization term into the training objective to encourage collaborative exploration of wide minima. However, like other measures, Inv. MV suffers from computational inefficiency. Calculating these measures typically requires either loss landscape sampling (as in $\epsilon$-sharpness, LPF, and Inv. MV) or computing second-order information such as the Hessian. In particular, Inv. MV involves locating boundary points via line search along worker directions, making it impractical for time-constrained training scenarios.
To address this limitation, we propose a relaxation of Inv. MV suitable for efficient, lightweight incorporation into the training objective. Instead of conducting the exhaustive boundary-point search, we directly treat current worker parameters as approximate boundary points, thus circumventing the costly line-search procedure altogether. This corresponds to adding the following regularization term to the objective in Equation 1, scaled by the regularization coefficient $\lambda_{r}$: $\lambda_{r}\mathcal{R}=-\frac{\lambda_{r}}{M}\sum_{i=1}^{M}\|x_{i}-x_{A}\|_{2}$, where $x_{A}$ denotes the average of the worker parameters. Essentially, this is equivalent to increasing the consensus distance, contrary to the typical practice in distributed training kong2021consensus , which aims to reduce it. We then derive the update that arises due to the presence of $\mathcal{R}$ (the full derivation steps can be found in Section E.1):
$-\lambda_{r}\frac{\partial\mathcal{R}}{\partial x_{m}}\overset{(\blacklozenge)}{=}\frac{\lambda_{r}}{M^{2}}\left(M\frac{x_{m}-x_{A}}{\|x_{m}-x_{A}\|}-\sum_{j=1}^{M}\frac{x_{j}-x_{A}}{\|x_{j}-x_{A}\|}\right)\overset{(\star)}{\approx}\lambda\frac{x_{m}-x_{A}}{\|x_{m}-x_{A}\|}.$ (2)
Observe that the expression on the right-hand side (RHS) of ()($\blacklozenge$)( ◆ ) contains two terms within parentheses. The first term exerts a force pushing the worker away from the average point $x_{A}$. The second term is the negative of the averaged normalized worker directions, which approaches zero when the workers are symmetrically distributed around $x_{A}$. Specifically, under the assumption that each normalized vector $\frac{x_{j}-x_{A}}{\|x_{j}-x_{A}\|}$ is uniformly distributed on the unit sphere in a $d$-dimensional space, its expectation is exactly zero. Thus, we omit this term in practice. Furthermore, absorbing the constant factor $M$ in the denominator into $\lambda_{r}$ ($\lambda=\frac{\lambda_{r}}{M}$) simplifies the expression, yielding the approximation on the RHS of ()($\star$)( ⋆ ) that exerts a unit-normed pushing force directly scaled with regularization coefficient $\lambda$.
$x_{m}\leftarrow x_{m}+(x_{A}-x_{m})\left(\alpha-\frac{\lambda}{\|x_{m}-x_{A}\|}\right)$ (3)
Recall that the distributed consensus step involves a pull force, implemented as the convex combination of the consensus variable $x_{C}$ and worker parameters $x_{m}$: $x_{m}\leftarrow(1-\alpha)x_{m}+\alpha x_{C}$. When our regularization term is active, it introduces a push force that opposes this pull step. Hence, we refer to our method as Distributed Pull-Push Force (DPPF). The push update is captured in Algorithm 1. Notice that when $x_{C}=x_{A}$, the pull and push updates elegantly combine into a single, concise expression presented in Equation 3. The term within parentheses on the right-hand side explicitly captures the interplay between the pull $(\alpha)$ and push $(\lambda)$ forces. This combined formulation is efficient, as it requires only a single step to apply both pull and push updates.
6 Theoretical Analysis
We begin by showing how the tug-of-war between DPPF’s pull ($\alpha$) and push ($\lambda$) forces determines the asymptotic valley width  [^2] We refer to a worker’s distance to $x_{A}$ as valley width since DPPF treats workers as valley boundaries., defined as the Euclidean distance between each worker and the global average $x_{A}$. Theorem 1 proves that, under mild assumptions, this width concentrates around the ratio $\lambda/\alpha$ when the learning rate decays and the number of workers is large.
Theorem 1 (DPPF: Final Valley Width).
Consider $M$ workers running DPPF with pull strength $\alpha$, push strength $\lambda$, communication period $\tau$, and local step size $\eta$. Let $x_{m,k}^{t}$ be worker $m$’s parameters at local iteration $t$ of communication round $k$, and $x_{A,k}$ the average over workers before the round. Define the post-update gap $\Delta_{m,k}^{+}:=x_{A,k}-x_{m,k}^{+}$. Assume unbiased stochastic gradients with bounded variance $\mathbb{E}[g_{m,k}^{t}]=\nabla f(x_{m,k}^{t}),\;\mathbb{E}\|g_{m,k}^{t}-\nabla f(x_{m,k}^{t})\|^{2}\leq\sigma_{0}^{2}\;$ for all $m,k,t$. Then
$\lim_{k\to\infty}\mathbb{E}\bigl{\|}\Delta_{m,k}^{+}\bigr{\|}=\tfrac{\lambda}{\alpha}+\mathcal{O}\!\bigl{(}\eta\sigma_{0}+M^{-1/2}\bigr{)},$
where the expectation is over stochastic randomness. Moreover, letting $\eta\to 0$ and $M\gg 1$ yields $\lim_{k\to\infty}\mathbb{E}\bigl{\|}\Delta_{m,k}^{+}\bigr{\|}={\lambda}/{\alpha}$.
Since $\lambda$ and $\alpha$ are user-chosen hyperparameters, we can preset the target valley size. In other words, DPPF embeds prior information about how wide a basin it should seek simply through the force ratio $\lambda/\alpha$. Next, we translate this controllable valley width into a PAC-Bayes guarantee.
Theorem 2 (DPPF: Wider Valley Tightens the Generalization Gap).
[…]
$\mathbb{E}_{x\sim Q_{r_{j}}}\![L_{\mathcal{D}}(x)]\leq\mathbb{E}_{x\sim Q_{r_{j}}}\![L_{S}(x)]+\underbrace{\sqrt{\frac{\,\tfrac{d}{2}(c_{r_{j}}-1-\log c_{r_{j}})+\frac{D_{0}}{2\sigma_{0}^{2}r_{j}^{1-\beta}}+\log\frac{nJ}{\delta}}{2(n-1)}}}_{\displaystyle\textstyle\text{gap}(r_{j})}.$
because $1-\beta>0$, $\text{gap}(r_{j+1})<\text{gap}(r_{j})$ for every consecutive pair in $\mathcal{G}$.
The bound’s leading term, $\text{gap}(r_{j})$, is strictly decreasing in $r_{j}$ because $1-\beta>0$. Hence, choosing a larger valley—attainable by increasing $\lambda/\alpha$—provably narrows the PAC-Bayes generalization gap under some technical conditions. Overall, Theorems 1 and 2 links a targeted valley knob (the push-to-pull ratio) to an explicit statistical benefit, a tighter generalization bound. We provide proofs of both theorems in Section E in the Appendix, which also includes insights into the arrangement of workers at the loss boundary and the non-convex convergence analysis.
7 Empirical Results
The empirical results presented here employ the increasing schedule for $\lambda$. An ablation study on different schedules is deferred to Section C.2 in the Appendix due to space constraints.
7.1 Push Mechanism with Soft-Consensus Methods
We integrate the proposed pull-push framework into existing soft-consensus distributed optimizers: EASGD, LSGD, and MGRAWA. These methods inherently apply a pulling force, guiding workers toward a consensus variable $x_{C}$. We also introduce SimpleAvg, which sets $x_{C}=x_{A}$, as a soft-consensus variant of LocalSGD stich2019local . We refer to SimpleAvg with the pull-push mechanism as DPPFSimpleAvg, and adopt the same naming convention when incorporating the push force into EASGD, LSGD, and MGRAWA. To see how much improvement in generalization we attain by introducing the wide minima seeking pushing force, we train ResNet-18 resnet models on CIFAR-10 cifar10 and CIFAR-100 cifar100 datasets using the vanilla distributed trainers and their DPPF variants. The experiments are conducted on 444 and 888 GPUs to assess the DPPF’s scalability. For the vanilla methods, we vary the pulling strength as follows: $\alpha\in\{0.05,0.1,0.3,0.5\}$, and for DPPF variants, $\alpha=0.1$ is fixed and the pushing force is varied as follows: $\lambda\in\{0.05,0.1,0.25,0.5,0.75\}$. In all the experiments, SGD is used as the underlying local optimizer. More details can be found in Section B.2 in the Appendix. The experiments are repeated for three different seeds and we report the test error as the average across the seeds, specifying the standard deviation in the subscript. (NC: not converged)
Table 2: Test performance of soft-consensus optimizers with and without DPPF. (NC: see Remark 3 in the Appendix).
CIFAR-10 CIFAR-100
4 Workers 8 Workers 4 Workers 8 Workers
SimpleAvg $4.24_{\pm 0.15}$ $4.31_{\pm 0.04}$ $21.19_{\pm 0.25}$ $21.44_{\pm 0.12}$
DPPFSimpleAvg $3.93_{\pm 0.09}$ $3.98_{\pm 0.10}$ $20.59_{\pm 0.06}$ $20.77_{\pm 0.20}$
EASGD $4.19_{\pm 0.21}$ $4.21_{\pm 0.17}$ $21.04_{\pm 0.22}$ $21.42_{\pm 0.06}$
DPPFEASGD $4.04_{\pm 0.08}$ $3.97_{\pm 0.11}$ $20.59_{\pm 0.29}$ $20.76_{\pm 0.17}$
LSGD $4.31_{\pm 0.03}$ $4.33_{\pm 0.13}$ $21.08_{\pm 0.34}$ $21.75_{\pm 0.43}$
DPPFLSGD NC NC NC NC
MGRAWA $4.35_{\pm 0.09}$ $4.22_{\pm 0.11}$ $21.31_{\pm 0.14}$ $21.04_{\pm 0.14}$
DPPFMGRAWA $4.03_{\pm 0.08}$ $3.99_{\pm 0.10}$ $20.46_{\pm 0.18}$ $20.87_{\pm 0.24}$
As shown in Table2, the push mechanism brings considerable improvements over the vanilla distributed optimizers that only have the pulling force. DPPF lowers test error by up to 0.3% on CIFAR-10 and 0.7% on CIFAR-100 experiments respectively. Among the DPPF variants that converged, we do not observe a significant difference in performance. Therefore, we use DPPFSimpleAvg moving forward to be able to execute pull and push updates in a single step, as mentioned in Equation 3 and refer to it simply as DPPF.
7.2 Comparison with Other Communication-Efficient Methods
Table 3: Comparison of DPPF with other communication-efficient distributed training methods.
Test Error (%)
RN-18
C10
PyNet
C100
RN-50
ImageNet
Comm. (%)
DDP SGD $4.33_{\pm 0.08}$ $19.10_{\pm 0.06}$ $23.83_{\pm 0.17}$ 100.0
LocalSGD $\tau=4$ $4.36_{\pm 0.06}$ $19.03_{\pm 0.28}$ $24.11_{\pm 0.04}$ 25.0
$\tau=8$ $4.40_{\pm 0.02}$ $18.82_{\pm 0.21}$ $24.21_{\pm 0.13}$ 12.5
$\tau=16$ $4.49_{\pm 0.21}$ $19.11_{\pm 0.21}$ $24.36_{\pm 0.19}$ 6.3
LocalSGD
+ QSR
$\tau_{\text{base}}=2$ $4.21_{\pm 0.08}$ $17.77_{\pm 0.06}$ $23.68_{\pm 0.02}$ 42.8
$\tau_{\text{base}}=4$ $4.32_{\pm 0.03}$ $18.06_{\pm 0.20}$ $23.49_{\pm 0.19}$ 16.1
$\tau_{\text{base}}=8$ $4.29_{\pm 0.04}$ $18.27_{\pm 0.26}$ $23.52_{\pm 0.21}$ 9.8
DPPF $\tau=4$ $\mathbf{3.93_{\pm 0.09}}$ $\mathbf{16.74_{\pm 0.03}}$ $\mathbf{23.07_{\pm 0.25}}$ 25.0
$\tau=8$ $\mathit{4.01_{\pm 0.05}}$ $\mathit{17.09_{\pm 0.12}}$ $\mathit{23.28_{\pm 0.22}}$ 12.5
$\tau=16$ $4.13_{\pm 0.08}$ $17.28_{\pm 0.04}$ $23.40_{\pm 0.15}$ 6.3
In this section, we compare DPPF with other communication-efficient methods, namely LocalSGD stich2019local and LocalSGD combined with QSR guquadratic , which, to the best of our knowledge, represents the current state of the art approach. As discussed previously, QSR proposes increasing the communication period $\tau$ throughout training to be inversely proportional to the squared learning rate guquadratic . Specifically, the authors schedule the communication period as $\tau_{t}=\max\left\{\tau_{\text{base}},\left\lfloor\left(\beta/\eta_{t}\right)^{2}\right\rfloor\right\},$ where $\tau_{\text{base}}$ is the minimum number of local steps taken before QSR is applied, and $\beta$ is a growth coefficient that controls how aggressively the communication period $\tau_{t}$ is updated based on changes in the learning rate $\eta_{t}$. In guquadratic , the authors experiment with $\tau_{\text{base}}\in\{2,4,8\}$ and $\beta\in\{0.2,0.25,0.3\}$, and keep $\eta_{\text{max}}=0.8$. To reproduce comparable effects in our own settings, we scale the $\beta$ values accordingly. Further training details are provided in Section B.3 of the Appendix.
Table 3 compares the test error and communication volumes of DPPF against LocalSGD and LocalSGD+QSR, across a diverse set of models (ResNet-18 resnet , PyramidNet(110,270) pyramidnet , and ResNet-50 resnet ) and datasets (CIFAR-10 cifar10 , CIFAR-100 cifar100 , and ImageNet imagenet ). We also quantify the communication volume as the ratio between the number of communication rounds required by each setting with respect to the local iteration count. We observe that DPPF consistently achieves the lowest test errors across all settings while operating at reduced communication costs. For instance, with $\tau=4$, DPPF attains a CIFAR-10 test error of 3.93%, outperforming all baselines—including DDP SGD (4.40%)—despite using only 25% of the communications. Similarly, on CIFAR-100 with PyramidNet, DPPF achieves 16.74% test error, again surpassing the rest, including the best LocalSGD+QSR setting, relying on 42.8% communication cost. On ImageNet with ResNet-50, DPPF achieves 23.07% test error at $\tau=4$, notably better than other communication-efficient methods, and DDP SGD (23.83%) while using a 4x lower than it. Importantly, DPPF also scales gracefully to longer communication intervals. With $\tau=8$ or $\tau=16$, it still outperforms most baselines while reducing communication volume to 12.5% and 6.3%, respectively. This highlights the robustness of our proposed pull-push mechanism in retaining generalization performance even under a tight communication budget, as also illustrated in Figure 1. These results collectively demonstrate that DPPF offers a gain on both generalization and communication efficiency.
7.3 DPPF Achieves SAM-like Performance While Being Communication-Efficient
Table 4: Test errors(%) reported by employing four different flatness encouraging schemes in distributed optimization.
DDP
SGD
DPPF
SGD
DDP
SAM
DPPF
SAM
4 Workers
C10
ResNet-18 $4.33_{\pm 0.08}$ $\mathbf{3.93_{\pm 0.09}}$ $3.97_{\pm 0.08}$ $\mathbf{3.74_{\pm 0.05}}$
WRN-16x8 $4.09_{\pm 0.10}$ $\mathbf{3.76_{\pm 0.06}}$ $3.72_{\pm 0.08}$ $\mathbf{3.70_{\pm 0.05}}$
PyNet(110,270) $3.94_{\pm 0.03}$ $\mathbf{3.11_{\pm 0.04}}$ $\mathbf{2.84_{\pm 0.10}}$ $2.89_{\pm 0.07}$
C100
ResNet-18 $21.29_{\pm 0.23}$ $\mathbf{20.59_{\pm 0.06}}$ $20.82_{\pm 0.14}$ $\mathbf{20.53_{\pm 0.13}}$
WRN-16x8 $20.10_{\pm 0.21}$ $\mathbf{18.99_{\pm 0.09}}$ $19.36_{\pm 0.06}$ $\mathbf{18.96_{\pm 0.18}}$
PyNet(110,270) $19.18_{\pm 0.10}$ $\mathbf{16.87_{\pm 0.18}}$ $16.50_{\pm 0.19}$ $\mathbf{15.68_{\pm 0.09}}$
8 Workers
C10
ResNet-18 $4.67_{\pm 0.17}$ $\mathbf{3.98_{\pm 0.10}}$ $4.23_{\pm 0.07}$ $\mathbf{3.92_{\pm 0.11}}$
WRN-16x8 $4.22_{\pm 0.08}$ $\mathbf{3.78_{\pm 0.12}}$ $3.79_{\pm 0.14}$ $\mathbf{3.79_{\pm 0.10}}$
PyNet(110,270) $4.08_{\pm 0.18}$ $\mathbf{3.18_{\pm 0.07}}$ $3.06_{\pm 0.05}$ $\mathbf{2.93_{\pm 0.07}}$
C100
ResNet-18 $21.26_{\pm 0.24}$ $\mathbf{20.77_{\pm 0.20}}$ $21.14_{\pm 0.14}$ $\mathbf{20.60_{\pm 0.12}}$
WRN-16x8 $20.58_{\pm 0.29}$ $\mathbf{18.90_{\pm 0.16}}$ $19.87_{\pm 0.15}$ $\mathbf{19.22_{\pm 0.13}}$
PyNet(110,270) $19.10_{\pm 0.06}$ $\mathbf{16.74_{\pm 0.03}}$ $16.68_{\pm 0.03}$ $\mathbf{15.94_{\pm 0.24}}$
In this section, we evaluate DPPF’s ability to seek flat minima at the distributed level by comparing it with SAM. As a baseline, we include DDP SGD, which uses standard gradient averaging with SGD as the local optimizer. Unlike SAM, which promotes flatness locally, DPPF enforces wide, high-quality minima through inter-worker repulsion. To compare local versus distributed flatness strategies, we include DDP SAM, which applies SAM locally with standard distributed gradient averaging. Finally, to evaluate the benefit of combining both local and distributed flatness-promoting mechanisms, we introduce DPPF SAM, which uses DPPF for distributed training and SAM as the local optimizer.
To compare these methods, we train three different models, namely ResNet-18 resnet , WideResNet-16x8 (WRN-16x8) wideresnet and PyramidNet-(110,270) (PyNet(110,270)) pyramidnet on CIFAR-10 (C10) cifar10 and CIFAR-100 (C100) cifar100 datasets. We run the experiments using 444 and 888 workers. Configurations with SGD optimizer are run for 400400400 epochs while those with SAM are run for 200200200 epochs for computational parity, as suggested in SAM . The SAM’s maximization hyperparameter is varied as $\rho\in\{0.05,0.1,0.2\}$. For DPPF, the communication period is fixed at $\tau=4$. More details can be found in Section B.4. Conclusions from Table 4 are consistent with that of Table 3 – our distributed flatness-promoting method, DPPF, outperforms standard DDP in generalization while retaining communication efficiency. Notably, DPPF SGD alone matches or surpasses DDP SAM in many cases, predominantly in CIFAR-100 experiments. Moreover, combining local and distributed flatness mechanisms (DPPF SAM) yields further improvements in generalization overall.
8 Ablation Studies
8.1 Analysis of the Pull-Push Mechanism
(a) Test error curves
(b) Consensus distance
Figure 2: Comparison of training with weaker pulling force and using the pull-push mechanism.
(c) Entire training + zoomed
Figure 3: Tug-of-war between pull and push forces.
Instead of introducing a pushing force to recover wide valleys, as done in DPPF, one might ask: why not simply reduce the pulling strength to keep workers apart? However, intuitive physical reasoning suggests the two approaches are not equivalent—having only a one-directional merging force inevitably leads to worker collapse. To test this, we conduct experiments using SimpleAvg without a pushing force and vary the pull strength across $\alpha=\{0.0001,0.005,0.01,0.05\}$. We compare these results with DPPF using $\alpha=0.1$ and $\lambda=0.5$, as reported in previous tables. All experiments are run with three random seeds. As shown in Figure 2(a), the wide-minima-seeking behavior enabled by DPPF’s pushing force cannot be replicated by merely weakening the pull. To further analyze this, we track the mean distance between workers and the average variable (consensus distance)—equivalently, the relaxed MV measure defined in Section 5—over the course of training (Figure 2(b)). The results show that, regardless of pulling strength, workers without a push force steadily collapse toward one another. This effect is consistent across SimpleAvg, EASGD, LSGD, and MGRAWA, and it restricts exploration of the loss landscape toward the end of training. We refer to this phenomenon as valley collapse, and our findings show that the presence of a push force is essential to preventing it. The valley collapse phenomenon is also visually evident in Figure 4(a).
Finally, we analyze the interplay between the pull and push forces, identifying the phases of training where the pulling force dominates and those where the pushing force prevails. We take one of the workers and highlight these phases as we plot the change of the worker’s distance to the average variable throughout the training. The results are shared in Figure 3. We see that the pulling force is stronger than the pushing force in the earlier phase of the training, yet the workers drift away from the average variable. In this phase, the exploration of the workers is propelled by the underlying local optimizer (SGD or variant), and the learning rate hasn’t decayed significantly yet to prevent the exploration. Close to convergence, the pushing force starts to overwhelm the pulling force, preventing the valley collapse and encouraging wider minima. At the end of the training, the interplay between the pull-push force is stabilized around 555, when $\lambda=0.5$ and $\alpha=0.1$, aligned with our Theorem 1.
Remark 1.
(Limitation of DPPF: Hyperparameters) Although DPPF has two tunable hyperparameters—the pull strength ($\alpha$) and the push strength ($\lambda$)—our ablation studies across various experimental settings show that the method consistently improves final performance over a wide range of $(\alpha,\lambda)$ values. Additional details are provided in Section D.2 of the Appendix.
8.2 Error Landscape Visualizations
[…]
(a) SimpleAvg
(b) DPPFSimpleAvg
Figure 4:  [^2] D Train Error (%) plots.
(c) SimpleAvg
(d) DPPFSimpleAvg
Figure 5:  [^3] D Train Error (%) plots.
Remark 2.
(Limitation of DPPF: Memory Usage) Similar to the soft-consensus methods, DPPF requires each worker to have an extra parameter vector for computing the average variable. Because this vector is needed only during communication rounds, it can be offloaded to the CPU between communications, easing GPU-memory pressure at the expense of minor I/O overhead.
9 Closing Remarks
In this work, we address the performance limitations of communication-efficient distributed training methods by proposing DPPF, a training strategy that introduces a push mechanism to prevent worker entrapment in sharp valleys of the loss landscape and to collaboratively recover flat minima. DPPF outperforms both the synchronous gradient-averaging method and other baselines by a considerable margin, while attaining comparable performance to SAM. Future work could apply the push mechanism to ensemble learning, and post-hoc refinement methods can further exploit the wide loss basins uncovered by DPPF. Additionally, extending DPPF to heterogeneous data settings and self-supervised learning presents promising research directions for general-purpose, scalable deep learning.
Appendix A Additional Related Works
Data Parallel Training of DNNs Local SGD has been shown to match the variance reduction properties and convergence rate of DDP SGD, achieving linear speedups with the number of workers stich2019local . In the conventional LocalSGD method, the communication period $\tau$ is set at the beginning of training and kept constant throughout. A series of theoretical works has shown that the convergence rate of LocalSGD is inversely proportional to the communication period $\tau$ whyandwhenlocalsgd ; stich2019local ; yu2019parallel , highlighting a trade-off between communication efficiency and model performance. This theoretical result has also been empirically validated in a study ortiz2021trade . Another notable work Gupta2020Stochastic proposes a hybrid approach and suggests performing All-Reduce SGD until a certain level of performance is reached and then switching to LocalSGD for final model averaging for improved performance. However, ortiz2021trade reports that the final performance highly depends on the transition point between the two training practices and usually, the optimal switching is late in the total training. Prior work has also focused on optimizing the communication period by incorporating adaptivity. kamp2014communication proposed a policy to skip or perform a consensus update in local communication based on the worker parameter variance. haddadpour2019local suggests linearly increasing the communication period as the training progresses. shen2021stl proposes a scheme in which the communication period is doubled whenever a two-fold learning rate drop is applied based on the predefined milestones. On the contrary, wang2019adaptive recommends starting with a high communication period and gradually decreasing toward the end of training. Although increasing the communication period usually hurts optimization, some recent works theoretically showed the benefits of increased communication periods in terms of generalization.
Flat Minima and Generalization The concept of seeking flat minima dates back to hochreiter1994simplifying . Since then, numerous studies have investigated the geometry of the loss landscapes of DNNs goodfellow2014qualitatively ; dauphin2014identifying ; baldassi2015subdominant ; li2018visualizing and the relationship between the flat minima and the generalization ability of the model keskar2016large ; jastrzkebski2017three ; andriushchenko2023modern . One early approach keskar2016large develops a sharpness metric that quantifies the maximum fluctuation in the loss value under controlled weight perturbations. The authors show that the minima from large-batch training are sharper and generalize worse than those from small-batch training. Another approach neyshabur2017exploring opposes the definition of the sharpness metric proposed in keskar2016large , criticizing it for falling short on accurately reflecting loss sharpness. They experiment with several norm-based sharpness measures that correlate well with generalization. Furthermore, jastrzkebski2017three characterizes how learning rate and batch size influence the final minima. A comprehensive study examining over 404040 complexity measures from theoretical bounds and empirical studies, along with their causal relationship to generalization, is provided in fantastic_measures . dinh2017sharp points out that one needs to be careful in designing measures to quantify sharpness for DNNs with ReLU activations, as simple reparameterizations can alter the geometry of the landscape. Finally, andriushchenko2023modern raises concerns about the validity of flatness measures in the modern DL.
Parameter Averaging Parameter averaging is also a widely adopted practice in Federated Learning (FL) mcmahan2017communication , where clients collaboratively train a DL model through a central server without sharing their local data due to privacy constraints. However, FL faces challenges arising from data heterogeneity and the limited participation of clients in global updates, which can degrade performance. Early approaches karimireddy2020scaffold ; acarfederated ; fan2022fedskip introduced statistical mechanisms to mitigate discrepancies between local updates and the global objective. More recent works  fedsam ; FedGAMMA have incorporated flat-minima-seeking optimization strategies into client objectives and locally estimated the sharpness of the global objective to improve generalization fedsmoo ; fedlesam . Beyond FL, model averaging is also leveraged in ensemble learning, where multiple models are trained independently on the full dataset while periodically sharing their weights to enhance robustness and performance jolicoeurpopulation ; fournier2024wash .
Appendix B Details of the Main Experiments
In this section, we offer further details on the primary experiments conducted, along with the best-performing hyperparameter configurations.
Licenses: Below we list the existing assets we use in our work and their corresponding license:
  • PyTorch Framework (version 2.6) - BSD-3-Clause
  • CIFAR-10/100 Datasets - Apache-2.0
  • ImageNet Dataset - Custom license
Computational Resources: We run experiments using both our local systems and cloud GPU services. Locally, we have 3 machines dedicated to the experiments conducted in this work and each machine is equipped with 4 x GTX-1080 GPUs and they are connected over Ethernet. We also reserve machines with 4 x H-100 GPUs from cloud services, mainly for the ImageNet experiments.
B.1 Generalization Gap vs. Sharpness Measures
[…]
Table 5: List of hyperparameters to be varied.
Hyperparameter Search Grid
Learning Rate [0.05,0.075,0.1][0.05,0.075,0.1][ 0.05 , 0.075 , 0.1 ]
Weight Decay [0,0.0001,0.01][0,0.0001,0.01][ 0 , 0.0001 , 0.01 ]
Momentum [0.1,0.5,0.9][0.1,0.5,0.9][ 0.1 , 0.5 , 0.9 ]
Batch Size [32,128,512][32,128,512][ 32 , 128 , 512 ]
Model Width [4,6,8][4,6,8][ 4 , 6 , 8 ]
The model we use is based on ResNet-18 architecture (aside from the varying convolutional channels) with skip connections and batchnorm layers. Additionally, since the model activation functions are ReLU, the model is scale-invariant dinh2017sharp , i.e. the model output can be preserved with re-parameterization of the model. Hence, we normalize all model components (convolution layers, linear layers, learnable batchnorm parameters) such that their Frobenius norms are made 111 before we conduct the sharpness analysis following bisla2022low . This normalization ensures that the conclusions drawn from the analysis cannot be altered by simple rescaling of intermediate model layers, thereby reflecting the true representativeness of the sharpness measures. To assess how well the sharpness measures are aligned with the generalization gap, we use the Kendall correlation coefficient that quantifies the similarity between the orderings of two lists. Below, we provide details on the measures that we inherit from the literature and our measure MV.
Shannon Entropy pereyra2017regularizing measures the confidence of a DNN based on the output distributions. Because having a confidence prediction signals that the model is overfitted, this sharpness measure can be expressed as negative of the Shannon Entropy which can be formulated as $\frac{1}{N}\sum_{i=1}^{n}\sum_{j=1}^{c}\phi_{j}(x;u_{i})\log\phi_{j}(x;u_{i})$ where $n$ is the number of samples, $u_{i}$ is the data sample $i$, $x$ is the model parameter vector, $\phi_{j}(x;u_{i})$ class $j$’s probability and $c$ is the number of classes. Also, note that this measure is not tied to the geometry of the landscape.
$\epsilon$-Sharpness keskar2016large can be expressed as the difference between the maximum loss obtained with perturbed model parameters and the original parameters where the perturbation is limited to a box with size $\epsilon$ for each parameter. The parameters that attain the maximum loss in the perturbed region can be determined by first calculating a full-batch gradient and mapping it to the perturbation space.
[…]
Low Pass Filter (LPF) bisla2022low accumulates the loss $M$ times due to Markov Chain Monte Carlo (MCMC) iterations for the convolution approximation. In each MCMC iteration, a random vector is drawn from Normal distribution, i.e. $\varepsilon\sim\mathcal{N}(0,\sigma I)$, and the loss is calculated at $\theta+\varepsilon$ contributes to the total sum which is then averaged across all MCMC iterations. In the experiments, we set $\sigma=0.01$ and $M=100$.
[…]
Inv. Mean Valley is the additive inverse of the Mean Valley metric which measures the valley width by pushing workers away from the average point. The measure is explained in detail in section˜4. We provide the pseudo-code that calculates this measure in algorithm˜2. Additionally, we conduct a study to assess the sensitivity of our measure on its hyperparameter $\kappa$. In particular, we vary $\kappa\in\{1.5,1.75,2.0,2.25,2.5,3.0,3.5,4.0\}$ and re-calculate the correlation strength when the analysis is carried out on the model parameters trained with and without data augmentation.
Table 6: Hyperparameter sensitivity analysis of the Inv. MV measure
$\mathbf{\kappa}$ 1.5 1.75 2.0 2.25 2.5 3.0 3.5 4.0
w/ Aug. 0.590 0.606 0.616 0.614 0.612 0.613 0.610 0.604
w/o Aug. 0.516 0.503 0.485 0.489 0.489 0.488 0.486 0.482
As can be seen from the results in Table 6, the Inv. MV Measure consistently maintains its ability to exhibit strong correlation across a wide range of $\kappa$ values.
Input : trained model parameters from $M$ workers $x_{1},x_{2},...,x_{M}$; threshold value $\kappa$ for valley boundary detection ($\kappa=2$ by default); step size $s$ for the resolution of the line search ($s=0.1$ by default).
Normalize all model parameters to make them scale-invariant.
$x_{A}\leftarrow 0$
for $m=1$ to $M$; do
 $x_{A}\leftarrow x_{A}+\frac{x_{m}}{M}$
end for
$l_{A}\leftarrow f(x_{A};D_{train})$
for $m=1$ to $M$; do
 $d_{m}\leftarrow\frac{x_{m}-x_{A}}{\|x_{m}-x_{A}\|_{2}}$
 $x_{m,b}\leftarrow x_{A}$
 while True do
    $x_{m,b}\leftarrow x_{m,b}+sd_{m}$
    $l_{m,b}\leftarrow f(x_{m,b};D_{train})$
    if $l_{m,b}\geq\kappa l_{A}$; then
         break
      end if
   end while
end for
$\Psi\leftarrow 0$
for $m=1$ to $M$; do
 $\Psi\leftarrow\Psi+\frac{\|x_{m,b}-x_{A}\|_{2}}{M}$
end for
Output: $-\Psi$
Algorithm 2 Inverse Mean Valley
B.2 DPPF with Soft-Consensus Methods
In the experiment set, we pair the wide-minima seeking pushing force, that results from incorporating Inv. MV regularization to the objective, with other Soft-Consensus Methods namely SimpleAvg, EASGD, LSGD, and MGRAWA. We refer to the paired versions with pushing force DPPFSimpleAvg, DPPFEASGD, DPPFLSGD and DPPFMGRAWA. As the underlying optimizer, we use SGD with a momentum value of 0.90.90.9 and, weight decay of $1e-3$. We train ResNet-18 models on CIFAR-10 and CIFAR-100 methods for 400400400 epochs. Per-GPU batch size is set to 128128128. For 4-GPU experiments, the learning rate is set to 0.10.10.1 and for 8-GPU experiments, the learning rate is scaled linearly with the total effective batch size hence it is 0.20.20.2. We only use the basic image augmentations on the train dataset, i.e. random horizontal flip and random cropping from padded images with a padding value of 444.
[…]
Table 7: Best $\lambda$ values of DPPF variants in different settings.
CIFAR-10 CIFAR-100
4 GPUs 8 GPUs 4 GPUs 8 GPUs
DPPFSimpleAvg 0.5 0.5 0.5 0.75
DPPFEASGD 0.75 0.75 0.5 0.5
DPPFMGRAWA 0.5 0.5 0.5 0.1
Remark 3.
[…]
B.3 Comparison with Communication-Efficient Methods
In the experiments with LocalSGD, we use fixed communication periods $\tau$ as specified in Table 3, and set $\alpha=1.0$ to achieve full consensus at each communication step.
For employing QSR on top of LocalSGD, we adopt the $\beta$ values reported in guquadratic . (Note: the original paper denotes this hyperparameter as $\alpha$; we use $\beta$ here to avoid confusion with the distributed pull strength.) The authors tuned $\beta\in\{0.2,0.25,0.3\}$ under a maximum learning rate of $\eta_{\text{max}}=0.8$. To preserve the same scheduling dynamics in our setting, we match the ratio $\frac{\beta}{\eta_{t}}$ in the QSR formulation:
$\tau_{t}=\max\left\{\tau_{\text{base}},\left\lfloor\left(\frac{\beta}{\eta_{t}}\right)^{2}\right\rfloor\right\}.$
Based on this, we scale the $\beta$ values proportionally to our chosen $\eta_{\text{max}}$ in each training setup. Specifically, we search:
  • $\beta\in\{0.025,0.03125,0.0375\}$ when $\eta_{\text{max}}=0.1$ in ResNet-18 training,
  • $\beta\in\{0.050,0.0625,0.075\}$ when $\eta_{\text{max}}=0.2$ in PyramidNet training,
  • $\beta\in\{0.1875,0.234375,0.28125\}$ when $\eta_{\text{max}}=0.75$ in ResNet-50 training.
We train ResNet-18, PyramidNet, and ResNet-50 models on the CIFAR-10, CIFAR-100, and ImageNet datasets using 4-GPU (GTX-1080), 8-GPU (GTX-1080), and 4-GPU (H100) setups, respectively. The batch sizes are set to 512 (CIFAR-10), 1024 (CIFAR-100), and 3072 (ImageNet). ResNet-18 and PyramidNet models are trained for 400 epochs, and ResNet-50 is trained for 200 epochs, following the training recipe in SAM . In the ImageNet experiments, we use a weight decay of 0.0001, while for CIFAR-10 and CIFAR-100, the weight decay is set to 0.001. To reduce computation, we tune the optimal $\beta$ for QSR on PyramidNet with CIFAR-100, and scale it proportionally with the learning rate in other settings.
For DPPF, we fix $\alpha=0.1$ and search over $\lambda\in\{0.5,0.75,1.0\}$ for CIFAR-10 and CIFAR-100 experiments, after seeing the trend in the first experiment set presented in Section B.2. For ImageNet experiments, we fix the ratio between push ($\lambda$) and pull ($\alpha$) forces to 10, i.e, $\lambda/\alpha=10$ and search over the following grid: $(\lambda,\alpha)\in\{(0.1,1.0),(0.5,5.0),(0.9,9.0)\}$. For LocalSGD and DPPF, we evaluate fixed communication periods $\tau\in\{4,8,16\}$. For LocalSGD+QSR, we search over $\tau\in\{2,4,8\}$, consistent with the experimental protocol in guquadratic . We repeat all the experiments for 333 seeds. We report the best hyperparameters in Table 8.
Table 8: Optimal hyperparameters reported for LocalSGD + QSR and DPPF
LocalSGD + QSR ($\beta$) DPPF ($\alpha,\lambda$)
$\tau_{\text{base}}=2$ $\tau_{\text{base}}=4$ $\tau_{\text{base}}=8$ $\tau=4$ $\tau=8$ $\tau=16$
ResNet-18 - C10 0.025 0.03125 0.03125 (0.1,0.5) (0.1,0.5) (0.1,0.5)
PyramidNet - C100 0.050 0.0625 0.0625 (0.1,1.0) (0.1,0.75) (0.1,0.75)
ResNet-50 - ImageNet 0.1875 0.234375 0.234375 (0.5, 5.0) (0.5, 5.0) (0.9, 9.0)
B.4 Comparison with DPP and SAM
Here we provide more details on the experiments that we compare DPPF with DDP, the most commonly adopted distributed training method. Note that we refer to DPPFSimpleAvg by DPPF as mentioned in 7.2. In this experiment set, we also switch the underlying optimizer from SGD to SAM which minimizes the maximum loss value attained in a region and it is known to encourage flatter, good-quality minima. These four combinations - DDP SGD, DPPF SGD, DDP SAM, and DPPF SAM — allow us to compare the effectiveness of flat-minima-seeking updates incorporated into the optimization at different levels. DDP SGD does not have any flat minima seeking mechanisms, DPPF SGD encourages wide minima discovery in the distributed level, DDP SAM promotes flatness with its local optimizer’s objective and DPPF SAM is equipped with both distributed and local level flat, wide minima encouraging mechanisms.
Table 9: Optimal $\lambda$ values of DPPF variants and optimal $\rho$ values of DDP SAM in various settings.
CIFAR-10 CIFAR-100
4 GPUs 8 GPUs 4 GPUs 8 GPUs
RN18 $\lambda$ DPPF SGD 0.5 0.5 0.5 0.75
DPPF SAM 0.05 0.05 0.1 0.15
$\rho$ DDP SAM 0.1 0.2 0.2 0.2
WRN
-16x8
$\lambda$ DPPF SGD 0.25 0.25 0.5 0.5
DPPF SAM 0.05 0.05 0.15 0.1
$\rho$ DDP SAM 0.1 0.2 0.2 0.2
PyNet $\lambda$ DPPF SGD 0.5 0.5 0.75 0.75
DPPF SAM 0.1 0.1 0.15 0.15
$\rho$ DDP SAM 0.1 0.2 0.2 0.2
We train ResNet-18, WideResNet-16x8 (WRN-16x8) and Pyramidnet-110 additive with alpha=270 (PyNet(110,270)) models on CIFAR-10 and CIFAR-100 datasets in 4-GPU and 8-GPU setups. For the methods with SGD, the models are trained for 400400400 epochs with a momentum value of 0.90.90.9 and a weight decay coefficient of 0.0010.0010.001. We want to note that most papers developing SAM variants, including the original SAM paper, report 0.00050.00050.0005 as the optimal weight decay factor for the SGD optimizer. In our experiments, we observe that setting SGD’s weight decay to 0.00050.00050.0005 undermines its performance and report 0.0010.0010.001 as its optimal weight decay value. The per-GPU batch size is fixed at 128128128 in all experiments and the learning rate is scaled linearly with the total effective batch size. Particularly, the learning rates are 0.10.10.1 and 0.20.20.2 for the 4-GPU and 8-GPU experiments respectively. For the settings with SAM, the models are trained for 200200200 epoch since a single iteration of SAM is equivalent to double SGD iterations due to gradient ascent and descent steps. SAM’s underlying optimizer is also SGD with the same hyperparameters specified previously. Again, we only use the basic image augmentations on the train dataset, i.e. random horizontal flip and random cropping from padded images with a padding value of 444.
For DDP SAM, we search SAM’s ascent step coefficient $\rho\in\{0.05,0.1,0.2\}$. Although SAM searches $\rho$ in a wider grid, only these $\rho$ values are reported as optimal in various settings. We initially also experimented with $\rho\in\{0.05,0.1,0.2,0.3,0.4\}$ with ResNet-18 and observed that higher values than $\rho=0.2$ starts hurting the performance on both CIFAR-10 and CIFAR-100. Hence we ultimately settle on $\rho\in\{0.05,0.1,0.2\}$.
For DPPF SGD, we use the same search space as before $\lambda\in\{0.05,0.1,0.25,0.5,0.75\}$. For DPPF SAM however, we fix $\rho=0.1$ and then opt for a more conservative search space as two wide-minima seeking mechanisms might overwhelm and hinder the minimization of classification loss. Particularly, we search $\lambda\in\{0.03,0.05,0.1,15,0.2\}$ in the DPPF SAM configuration. We also fix the pulling force $\alpha=0.1$. Ideally, all hyperparameters should be searched jointly, but this would exponentially increase the number of experiments. Therefore, we focus on varying $\lambda$ alone which is the pushing force. We believe the generalization performance of this setting could be further improved with a more thorough, comprehensive hyperparameter search that also includes varying $\rho$ and $\alpha$. Similar to the previous experiments set, we run each configuration for 3 different seeds 182,437,965182,437,965182 , 437 , 965 and each configuration’s performance is determined by averaging the statistics across seeds. In Table 9, we share the hyperparameters of the best configurations for reproducibility.
Appendix C Details of the Ablation Studies
Here, we give details on the ablation studies presented in the main body of the paper. In all ablation studies, we run training or conduct analysis in a 4-GPU training setup unless otherwise stated.
C.1 Pull-Push Mechanism
Here we give details on the experiments to analyze the importance of pushing mechanisms and the interplay between the pull-push forces throughout the training. First, to demonstrate the necessity of our pushing force, we compare DPPFSimpleAvg that has pulling force $\alpha=0.1$ and pushing force $\lambda=0.5$ with vanilla SimpleAvg methods that have weaker pulling forces, particularly $\alpha\in\{0.001,0.005,0.01,0.05\}$. We train ResNet-18 models on the CIFAR-100 dataset in these configurations for 333 seeds 182,437,965182,437,965182 , 437 , 965 and average test errors curves across seeds to obtain the plot in 2(a) where the shaded regions indicate the standard deviation. We also log the Euclidian distance workers to the average variable during the training at every iteration, i.e. $\|x_{m}-x_{A}\|_{2}$ for each worker $m$. For 2(b), we calculate the simplified MV, which is the average of these distances and plot its change with respect to training iterations.
In addition to logging the distance between each worker, we also record the strength of the applied pulling force and the pushing force. At each iteration, we compare the strength of these two forces to characterize the interplay between them. The results are presented in 3.
C.2 How to Schedule $\lambda$?
We train ResNet-18 models on the CIFAR-100 dataset to compare the effect of different schedulings of the pushing force $\lambda$ to the end performance. Particularly, we compare three schedulings: fixed, decreasing and increasing throughout the training. These schedulings are plotted in 6(c) when $\lambda=0.5$ and below, we provide more details to each scheduling:
  • Fixed: The strength of the pushing force $\lambda$ is kept constant throughout the training.
  • Decreasing: The $\lambda$ value is decayed in parallel with the learning rate. Since we are using cosine annealing scheduler for the learning rate, at any iteration $t$, $\lambda_{t}=\frac{\lambda}{2}\left(1+\cos\left(\frac{t}{T}\pi\right)\right)$ where $T$ is the total number of iterations.
  • Increasing: In this setting, the strength of $\lambda$ is amplified towards the end of the training. We again base the amplification on the learning rate for simplicity and use flipped cosine annealing for scheduling. More specifically, $\lambda_{t}=\frac{\lambda}{2}\left(1-\cos\left(\frac{t}{T}\pi\right)\right)$ where $T$ is the total number of iterations.
(a) Test errors (%)
(b) Consensus distance
(c) Change of $\lambda$
Figure 6: Comparison of different schedulings.
For each scheduling, we repeat the experiment for 555 different seeds 42,182,437,965,128342,182,437,965,128342 , 182 , 437 , 965 , 1283 and compare the test error of different scheduling averaged over seeds. The results in 6(a) show that increasing the strength of $\lambda$ towards the end of training is more effective, as recovering wide basins becomes more important at the end. More specifically, the test errors and standard deviations, calculated across 555 seeds, are $20.84_{\pm 0.41}$, $21.05_{\pm 0.17}$, $20.61_{\pm 0.15}$ for fixed, increasing, and decreasing schedules, respectively.
Appendix D Additional Results and Ablation Studies
D.1 Ablation Study on the Second Term
In Section 5, we present the update rule that arises from the Simplified MV ($R$) term in the objective. However, in practice we execute the simplified update rule by only keeping the first term. Let us consider the full, original update with both terms. The full update expression is as follows (as proven in Section E.1):
$\frac{\partial R}{\partial x_{m}}=-\frac{\lambda}{M^{2}}\left(M\frac{d_{m}}{\|d_{m}\|}-\sum_{j=1}^{M}\frac{d_{j}}{\|d_{j}\|}\right).$
where $d_{m}=x_{m}-x_{A}$. In practice, we drop the second term in the parentheses because when the workers are symmetrically spread around the average variable, its value is close to 0. In such a case, the overall update can be approximated as:
$-\frac{\lambda}{M^{2}}\left(M\frac{d_{m}}{\|d_{m}\|}-\sum_{j=1}^{M}\frac{d_{j}}{\|d_{j}\|}\right)\approx-\frac{\lambda}{M}\frac{d_{m}}{\|d_{m}\|}$
We also empirically check if this is the case. Let $T_{1}=-\frac{\lambda}{M}\frac{d_{m}}{\|d_{m}\|}$ and $T_{2}=\frac{\lambda}{M^{2}}\sum_{j=1}^{M}\frac{d_{j}}{\|d_{j}\|}$ hence the overall update can be expressed as $T_{1}+T_{2}$. We run the experiment with $M=4$ workers and plot how the Euclidean norms of $T_{1}$, $T_{2}$, and $T_{1}+T_{2}$ change during the training. To check the validity of our claim empirically, we also scale the norm of $T_{1}$. The results in Figure 7 reveal that indeed the simplified expression $-\lambda\frac{1}{M}\frac{d_{m}}{\|d_{m}\|}$ is a good proxy to the actual update. Although some fluctuations are not captured, as perfect symmetry of workers around the average variable is not always ensured, we did not observe any change in the final performance. Besides, the simplified update is more communication-efficient as the calculation of the second term requires either an additional communication round among the workers, or a costlier communication to retrieve the model copies from all the workers.
Figure 7: Ablation on the presence of the second term
D.2 Ablation on DPPF’s Hyperparameter Sensitivity
In this section, we investigate how sensitive the DPPF is to the hyperparameter selections. Particularly, DPPF has two hyperparameters: the pull $\alpha$ and the push $\lambda$ force strengths. The theoretical analysis reveals that the final valley width found by DPPF is governed by the ratio of the push and pull forces, i.e., $\lambda/\alpha$. We consider training the PyramidNet(270,110) on CIFAR-100 for 400 epochs, a model with more than enough capacity that can potentially suffer from overfitting. In our first sensitivity analysis, we fix $\alpha=0.5$ and try different values of $\lambda$ for DPPF training. In particular, we use $\lambda\in\{0.1,0.25,0.5,1.0,2.5,5,7.5,10\}$ so that the DPPF finds valleys with different width. We share the final test errors (averaged across 3 seeds) and the valley width side-by-side in Figure 8.
(a) Test errors (%)
(b) Valley width
Figure 8: PyramidNet training on CIFAR-100 dataset with 4 workers.
[…]
(a) Average model norm
(b) Valley width over norm
Figure 9: PyramidNet training on CIFAR-100 dataset with 4 workers.
Figure 9(a) illustrates how the norm of the average variable ($|x_{A}|2$) evolves throughout training when different push force strengths are applied. We observe that stronger push forces lead to higher final norms of the average variable. In Figure 9(b), we report the ratio between the valley width, defined as $\frac{1}{M}\sum{i=1}^{M}|x_{i}-x_{A}|_{2}$, and the norm of the average variable $|x_{A}|_{2}$, measured at the end of training. This ratio grows approximately exponentially with increasing $\lambda$, and appears to diverge for large $\lambda$ values.
Recall from Figure 8(a) that the generalization improvements brought by the push force begin to saturate or degrade beyond $\lambda=5$. This suggests the existence of a critical $\lambda$—and potentially a critical valley-width-to-norm ratio—up to which consistent generalization benefits are observed. While DPPF yields substantial generalization improvements over a wide range of $\lambda$ values (for fixed $\alpha$), future work could aim to identify this critical threshold, potentially enabling a hyperparameter-free variant of DPPF. One promising direction could involve estimating a lower bound on the Lipschitz constant of the DNN for a given dataset, which may provide guidance on appropriate values for $\lambda$.
Previously, we examined how the performance improvements of DPPF vary with different values of $\lambda$ while keeping the pull force fixed at $\alpha=0.5$. Now, we fix the final valley width by preserving the ratio $\lambda/\alpha$, and explore different $(\lambda,\alpha)$ pairs to investigate how test error is affected. Before presenting our results, we report the performance achieved by the standard gradient-averaging scheme (DDP-SGD) across various experimental settings as a reference, as shown in Table 10.
Table 10: Test errors attained by DDP SGD Training
ResNet-18
CIFAR-10
4 Workers
400 Epochs
PyramidNet
CIFAR-100
4 Workers
400 Epochs
ResNet-50
ImageNet
4 Workers
200 Epochs
$4.33_{\pm 0.08}$ $19.18_{\pm 0.10}$ $23.83_{\pm 0.17}$
Unless otherwise stated, we maintain consistent experimental settings throughout our analysis. For CIFAR-10 and CIFAR-100 experiments, we vary the communication period $\tau\in\{4,8,16\}$ and use $(\lambda,\alpha)\in\{(0.1,0.5),(0.5,2.5),(0.9,4.5)\}$ to ensure that the final valley width—governed by the ratio $\lambda/\alpha$—remains fixed at 555. Additionally, we include experiments with a shorter training schedule of 200 epochs, in contrast to the default 400 epochs. For ImageNet experiments, we use $(\lambda,\alpha)\in\{(0.1,1.0),(0.5,5.0),(0.9,9.0)\}$ to keep the valley size fixed at 101010 and only provide 200 epoch training, the default recipe. Figures 10, 11, and 12 present the hyperparameter sensitivity analysis in the form of heatmaps.
(a) Training for 200 epochs
(b) Training for 400 epochs
Figure 10: ResNet-18 training on CIFAR-10 dataset with 4 workers.
(a) Training for 200 epochs
(b) Training for 400 epochs
Figure 11: PyramidNet training on CIFAR-100 dataset with 4 workers.
ResNet-18, CIFAR-10: In Figure 10, we observe that DPPF maintains a consistent level of performance across different communication periods and $(\lambda,\alpha)$ pairs. However, some trends emerge when comparing training durations: with shorter training (200 epochs), configurations with larger $\alpha$ values tend to yield better test error. In contrast, for longer training (400 epochs), intermediate values of $\alpha$ (e.g., (0.5,2.5)(0.5,2.5)( 0.5 , 2.5 )) perform slightly better, suggesting that the optimal balance between push and pull may shift depending on the training duration.
PyramidNet, CIFAR-100: In Figure 11, we observe that shorter training with 200 epochs does not have a clear trend, whereas the final performance remains stable across different configurations. More importantly, even with only 200 epochs, DPPF consistently outperforms 400-epoch DDP training across all combinations of communication period and pull-push strengths by a significant margin. In the 400-epoch setting, the final performance remains relatively stable across different hyperparameter choices, with the exception of a few outliers under $\tau=16$. This favorable reduction in sensitivity is likely due to the longer training duration and the over-parameterized nature of the PyramidNet model for CIFAR-100 data, which together enable robust convergence across a broader range of configurations.
Figure 12: ResNet-50 training on ImageNet dataset with 4 workers
ResNet-50, ImageNet: In Figure 12, a clear trend emerges: this setting benefits from more frequent communication and stronger pull forces (i.e., larger $\alpha$). This suggests that tighter synchronization helps stabilize training for large-scale datasets like ImageNet, possibly due to the increased complexity and depth of ResNet-50. Despite this, 8 out of the 9 DPPF configurations either match or significantly outperform the baseline DDP-SGD in terms of test error, all while offering improved communication efficiency. These results highlight the robustness and practical value of DPPF, even under varied hyperparameter settings. It would be valuable to extend this analysis to larger models on ImageNet, such as ResNet-101, ResNet-152, or Vision Transformers dosovitskiy2020image , to assess whether the observed sensitivity trends and performance gains scale with model size and architecture.
D.3 Is DPPF Essentially an On-the-Fly SWA?
Due to the arrangement of workers in the loss landscape driven by DPPF’s pushing force, one might wonder if DPPF is emulating Stochastic Weight Averaging (SWA) during training. In SWA izmailovaveraging , it has been observed that the minima found by SGD often lie at the edges of the valley, rather than at its center. To address this, the authors propose continuing training after convergence, using a cyclical learning rate to explore different edges of the valley. The SGD solutions obtained are then averaged to locate the center of the valley (SWA solution) where more robust, better solutions lie izmailovaveraging . Since all edge solutions and the SWA solution lie in the same basin, the trajectory between any of the two solutions does not encounter any loss barriers which is qualitatively verified in izmailovaveraging .
To address the question of whether DPPFSimpleAVG acts as an on-the-fly SWA for SimpleAVG, we check whether the last observation holds in our case. For this purpose, we plot how the loss and error (%) along the trajectory between SimpleAVG and DPPFSimpleAVG changes. Let $x_{sa}$ denote the solution from SimpleAVG and $x_{dppf}$ the solution obtained by DPPFSimpleAVG. We take a convex combination among these two solutions which can be expressed as $x_{C}=\alpha x_{dppf}+(1-\alpha)x_{sa}$ for $\alpha\in[0,1]$. For all $x_{C}$, we record the training loss, training error, test loss, and test error. The plots in fig.˜13 show that there is a large loss barrier between the solutions of SimpleAVG and DPPFSimpleAVG. This implies that DPPF does not simply emulate SWA, and it encourages the recovery of separate, wider valleys with good-quality solutions.
(a) $\alpha$ vs. loss
(b) $\alpha$ vs. error (%)
Figure 13: Change in training loss, training error, test loss, and test error along the trajectory between SimpleAVG and DPPFSimpleAVG solutions.
Appendix E Full Proofs of Theoretical Analysis
E.1 Derivation of the Update Rule by Minimizing the Relaxed Inv. MV Term
Recall that the relaxed Inv. MV measure is mathematically expressed as follows:
$R=-\frac{1}{M}\sum_{i=1}^{M}\|x_{i}-x_{A}\|_{2}\hskip 5.69054pt\text{where}\hskip 5.69054ptx_{A}=\frac{1}{M}\sum_{i=1}^{M}x_{i}$
Now, we want to derive the gradient of the term with respect to worker $m$, i.e. $\frac{\partial R}{\partial x_{m}}$. Let us also write $d_{i}=x_{i}-x_{A}$ so we can also write:
$R=-\frac{1}{M}\sum_{i=1}^{M}\sqrt{d_{i}^{T}d_{i}}\hskip 5.69054pt\text{where}\hskip 5.69054ptd_{i}=x_{i}-\frac{1}{M}\sum_{j=1}^{M}x_{j}$
We have two different cases to consider to find $\frac{\partial R}{\partial x_{m}}$. This is because the calculation of $x_{A}$ includes all the model parameters hence, $x_{m}$ is present in all terms of $R$’s summation due to $x_{A}$. Besides, $i=m$ includes the copy of $x_{m}$ itself inside the norm. These two cases require separate treatment to derive the update rule accurately.
1) $i=m$: We start by using the chain rule:
$\frac{\partial R}{\partial x_{i}}=\frac{\partial R}{\partial\sqrt{d_{i}^{T}d_{i}}}\frac{\partial\sqrt{d_{i}^{T}d_{i}}}{\partial d_{i}}\frac{\partial d_{i}}{\partial x_{i}}$
The individual terms in the chain rule are equal to the following:
$\frac{\partial R}{\partial\sqrt{d_{i}^{T}d_{i}}}=\frac{-1}{M},\hskip 5.69054pt\frac{\partial\sqrt{d_{i}^{T}d_{i}}}{\partial d_{i}}=\frac{d_{i}}{\sqrt{d_{i}^{T}d_{i}}},\hskip 5.69054pt\frac{\partial d_{i}}{\partial x_{i}}=\frac{M-1}{M}$
Plugging in the expression of the individual partial derivatives in the chain rule and re-writing $d_{i}=x_{i}-x_{A}$ we obtain the following:
$\frac{\partial R}{\partial\sqrt{d_{i}^{T}d_{i}}}\frac{\partial\sqrt{d_{i}^{T}d_{i}}}{\partial d_{i}}\frac{\partial d_{i}}{\partial x_{m}}=-\frac{M-1}{M^{2}}\frac{x_{m}-x_{A}}{\|x_{m}-x_{A}\|_{2}}$ (4)
This concludes the part of the update that arises from the case $m=i$.
2) $i\neq m$: We again start by using the chain rule:
$\frac{\partial R}{\partial x_{i}}=\frac{\partial R}{\partial\sqrt{d_{i}^{T}d_{i}}}\frac{\partial\sqrt{d_{i}^{T}d_{i}}}{\partial d_{i}}\frac{\partial d_{i}}{\partial x_{i}}$
Again we separately express the partial derivatives in the chain rule:
$\frac{\partial R}{\partial\sqrt{d_{i}^{T}d_{i}}}=\frac{-1}{M},\hskip 5.69054pt\frac{\partial\sqrt{d_{i}^{T}d_{i}}}{\partial d_{i}}=\frac{d_{i}}{\sqrt{d_{i}^{T}d_{i}}},\hskip 5.69054pt\frac{\partial d_{i}}{\partial x_{i}}=\frac{-1}{M}$
Notice that this time $\frac{\partial d_{i}}{\partial x_{i}}$ is different since $x_{m}$ only participates in $d_{i}$ due to the presence of $x_{A}$. Putting everything together we get the following for all $i\neq m$:
$\frac{\partial R}{\partial\sqrt{d_{i}^{T}d_{i}}}\frac{\partial\sqrt{d_{i}^{T}d_{i}}}{\partial d_{i}}\frac{\partial d_{i}}{\partial x_{m}}=\frac{1}{M^{2}}\frac{x_{i}-x_{A}}{\|x_{i}-x_{A}\|_{2}}$ (5)
Finally, by combining 4 and 5, we obtain the following overall update:
$\frac{\partial R}{\partial x_{m}}=-\frac{1}{M^{2}}\left((M-1)\frac{d_{m}}{\|d_{m}\|}-\sum_{j\neq m}^{M}\frac{d_{j}}{\|d_{j}\|}\right)=-\frac{1}{M^{2}}\left(M\frac{d_{m}}{\|d_{m}\|}-\sum_{j=1}^{M}\frac{d_{j}}{\|d_{j}\|}\right).$
This concludes the derivation.
E.2 Valley Width and Generalization Guarantees
Consider a distributed training of a DNN performed with $M$ workers by optimizing the loss function $f$. We define the following notations for communication round $k$ and worker $m$:
[…]
$\Delta_{m,k}$ $=x_{A,k}-x_{m,k}\quad\text{gap vector before distributed update},$
$\Delta_{m,k}^{+}$ $=x_{A,k}^{+}-x_{m,k}^{+}\quad\text{gap vector after distributed update},$
$r_{m,k}$ $=\lVert\Delta_{m,k}\rVert,\quad u_{m,k}=\Delta_{m,k}/r_{m,k}\quad(\lVert u_{m,k}\rVert=1),$
$C$ $=1-\alpha\quad(0<C<1).$
We also assume that the workers observe independent and unbiased gradients ($g$’s) with bounded variance during the training:
$\mathbb{E}\|g_{m,k}^{t}-\nabla f(x_{m,k}^{t})\|^{2}\leq\sigma_{0}^{2}\;\;\;\text{for all}\;\;\;m,k,t.$
Lemma 1.
Let $G_{j,k}=\sum_{t=1}^{\tau}g_{j,k}^{t}$ be the sum of local gradients calculated by worker $j$ between communication rounds $k$ and $k+1$. Also, $\bar{G}_{k}=\frac{1}{M}\sum_{i=1}^{M}G_{i,k}$ denotes the average of the local gradients across workers. Based on the bounded variance assumption of the gradients, we have:
$\mathbb{E}\bigl{[}\|G_{m,k}-\bar{G}_{k}\|\bigr{]}\leq\sqrt{\tfrac{M+1}{M}}\sqrt{\tau}\sigma_{0}$
Proof.
Using the definition that $G_{m,k}=\sum_{t=1}^{\tau}g_{m,k}^{t}$, we write the following:
$\bar{G}_{k}=\frac{1}{M}\sum_{j=1}^{M}G_{j,k},\qquad\mathbb{E}\bigl{[}\|\bar{G}_{k}\|^{2}\bigr{]}=\frac{1}{M^{2}}\sum_{j=1}^{M}\mathbb{E}\bigl{[}\|G_{j,k}\|^{2}\bigr{]}=\frac{\tau\sigma_{0}^{2}}{M}.$
$\mathbb{E}\bigl{[}\|G_{m,k}-\bar{G}_{k}\|^{2}\bigr{]}$ $=\mathbb{E}\bigl{[}\|G_{m,k}\|^{2}\bigr{]}+\mathbb{E}\bigl{[}\|\bar{G}_{k}\|^{2}\bigr{]}-2\mathbb{E}\langle G_{m,k},\bar{G}_{k}\rangle$
$=\tau\sigma_{0}^{2}+\frac{\tau\sigma_{0}^{2}}{M}\quad(\text{cross term }=0)$
$=\tau\sigma_{0}^{2}\!\Bigl{(}1+\frac{1}{M}\Bigr{)}.$
Finally, by using the Cauchy-Schwarz, we can bound the first moment as:
$\mathbb{E}\bigl{[}\|G_{m,k}-\bar{G}_{k}\|\bigr{]}\leq\sqrt{\mathbb{E}\bigl{[}\|G_{m,k}-\bar{G}_{k}\|^{2}\bigr{]}}=\sqrt{\tfrac{M+1}{M}}\sqrt{\tau}\sigma_{0}$
Lemma 2.
Let us assume that $u_{i}\in\mathbb{R}^{d}$ is a column vector drawn independently and uniformly from the unit sphere $\mathbb{S}^{d-1}$ so that the norms are 111 surely. For $\bar{u}_{k}=\frac{1}{M}\sum_{i=1}^{M}u_{i,k}$, we can write the following upper bound:
$\mathbb{E}\left[\left\|\bar{u}_{k}\right\|_{2}\right]\leq\frac{1}{\sqrt{M}}$
Proof.
We start by writing:
$\mathbb{E}\left[\left\|\bar{u}_{k}\right\|_{2}^{2}\right]$ $=\mathbb{E}\left[\bar{u}_{k}^{\top}\bar{u}_{k}\right]$
$=\frac{1}{M^{2}}\sum_{i,j=1}^{M}\mathbb{E}\left[u_{i,k}^{\top}u_{j,k}\right]$
$=\frac{1}{M^{2}}\left(\sum_{i=1}^{M}\mathbb{E}\left[\left\|u_{i,k}\right\|_{2}^{2}\right]+\sum_{i\neq j}^{M}\mathbb{E}\left[u_{i,k}^{\top}u_{j,k}\right]\right).$
Since $\mathbb{E}\left[\left\|u_{i,k}\right\|_{2}^{2}\right]=1$ by construction and the cross terms are 0 due to $\mathbb{E}[u_{i,k}]=\mathbf{0}$ and independence, we arrive at $\mathbb{E}\left[\left\|\bar{u}_{k}\right\|_{2}^{2}\right]=\frac{1}{M}$. Then, using Cauchy-Schwarz, we get:
$\mathbb{E}\left[\left\|\bar{u}_{k}\right\|_{2}\right]\leq\sqrt{\mathbb{E}\left[\left\|\bar{u}_{k}\right\|_{2}^{2}\right]}=\frac{1}{\sqrt{M}}$
Theorem 3 (Valley Width under DPPF).
Consider a distributed training of a DNN performed by $M$ workers using DPPF with pull and push force strengths $\alpha$ and $\lambda$ respectively. The communication period is $\tau$, and the workers locally run SGD with learning rate $\eta$. Denote the worker index, communication round, and local iteration with $m,k$, and $t$ respectively. We assume bounded variance: $\mathbb{E}\|g_{m,k}^{t}-\nabla f(x_{m,k}^{t})\|^{2}\leq\sigma_{0}^{2}\;\;\text{for all}\;\;m,k,t$. Now, define the gap vector $\Delta_{m,k}^{+}=x_{A,k}-x_{m,k}^{+}$ as the distance between worker $m$ and the worker average after the communication update $k$. Asymptotically, we obtain:
$\boxed{\lim_{k\to\infty}\mathbb{E}\bigl{[}\lVert\Delta_{m,k}^{+}\rVert\bigr{]}=\frac{\lambda}{\alpha}+\mathcal{O}\!\Bigl{(}\eta\sigma_{0}+\tfrac{1}{\sqrt{M}}\Bigr{)}}$
Furthermore, with $\eta\to 0$ and many workers $M\gg 1$, we have $\lim_{k\to\infty}\mathbb{E}\bigl{[}\lVert\Delta_{m,k}^{+}\rVert\bigr{]}=\frac{\lambda}{\alpha}.$
Proof.
We start by expressing how the gap changes after the distributed pull-push update.
$x_{m,k}^{+}$ $=x_{m,k}+\Delta_{m,k}\left(\alpha-\lambda\frac{1}{\|\Delta_{m,k}\|}\right)$ (6)
$x_{A,k}^{+}$ $=\frac{1}{M}\sum_{i=1}^{M}x_{i,k}^{+}=\underbrace{\frac{1}{M}\sum_{i=1}^{M}x_{i,k}}_{x_{A,k}}+\alpha\underbrace{\frac{1}{M}\sum_{i=1}^{M}\Delta_{i,k}}_{0}-\frac{\lambda}{M}\sum_{i=1}^{M}u_{i,k}$ (7)
By subtracting Equation 6 from Equation 7, we can write:
$\Delta_{m,k}^{+}=(1-\alpha)\Delta_{m,k}+\lambda u_{m,k}-\frac{\lambda}{M}\sum_{i=1}^{M}u_{i,k}$ (8)
However, we want to extract the recurrence between how the post-update gap changes. To this end, we need to relate $\Delta_{m,k-1}^{+}$ term to $\Delta_{m,k}^{+}$. Let us consider the local gradient steps between communication rounds $k-1$ and $k$.
$x_{m,k}=x_{m,k-1}^{+}-\eta G_{m,k-1}\;\;\;\text{where}\;\;\;G_{m,k-1}=\sum_{t=1}^{\tau}g_{m,k-1}^{t}$ (9)
Here $g_{m,k-1}^{t}$ is the gradient update that worker $m$ takes at $t^{\text{th}}$ local iteration after communication round $k-1$. And we use $G_{m,k-1}$ to denote the cumulative local updates. Let us also express:
$x_{A,k}=\frac{1}{M}\sum_{i=1}^{M}x_{i,k}=\underbrace{\frac{1}{M}\sum_{i=1}^{M}x_{i,k-1}^{+}}_{x_{A,k-1}^{+}}-\frac{\eta}{M}\sum_{i=1}^{M}G_{i,k-1}$ (10)
Let us all define $\bar{G}_{k-1}=\frac{1}{M}\sum_{i=1}^{M}G_{i,k-1}$ to denote the average of the local gradients across workers. By subtracting Equation 9 from Equation 10, we arrive at:
$x_{A,k}-x_{m,k}$ $=x_{A,k-1}^{+}-x_{m,k-1}^{+}-\eta\left(\bar{G}_{k-1}-G_{m,k-1}\right)$ (11)
$\Delta_{m,k}$ $=\Delta_{m,k-1}^{+}-\eta Z_{m,k-1}\;\;\;\text{where}\;\;\;Z_{m,k-1}=\bar{G}_{k-1}-G_{m,k-1}$ (12)
By plugging Equation 12 into Equation 8, we can obtain the following recurrence relation:
$\Delta_{m,k}^{+}=(1-\alpha)\Delta_{m,k-1}^{+}-\eta(1-\alpha)Z_{m,k-1}+\lambda u_{m,k}-\lambda\bar{u}_{k}$ (13)
where we expressed $\frac{1}{M}\sum_{i=1}^{M}u_{i,k}$ as $\bar{u}_{k}$. We now start from the recurrence:
$\Delta_{m,k}^{+}=C\Delta_{m,k-1}^{+}-\eta CZ_{m,k-1}+\lambda u_{m,k}-\lambda\bar{u}_{k}$ (14)
Taking norms and applying the triangle inequality:
$\|\Delta_{m,k}^{+}\|$ $\leq C\|\Delta_{m,k-1}^{+}\|+\eta C\|Z_{m,k-1}\|+\lambda\|u_{m,k}-\bar{u}_{k}\|$ (15)
$\leq C\|\Delta_{m,k-1}^{+}\|+\eta C\|Z_{m,k-1}\|+\lambda(1+\|\bar{u}_{k}\|)$ (16)
where we used that $\|u_{m,k}\|=1$ by construction. For ease of notation, let us define $r_{k+1}=\mathbb{E}\!\bigl{[}\lVert\Delta_{m,k+1}^{+}\rVert\bigr{]}$. We express $\mathbb{E}Z_{m,k}$ using Lemma 1. Additionally, we treat $u_{i}\in\mathbb{R}^{d}$ as a column vector drawn uniformly from the unit sphere $\mathbb{S}^{d-1}$ so that the norms are 111 surely. Also, $u_{i}$’s are drawn independently, a valid assumption as the independent stochastic noise determines the location of $x_{i,k}$. By invoking Lemma 2 on $\|\bar{u}_{k}\|$, we write the following:
$r_{k+1}$ $=\mathbb{E}\bigl{[}\lVert\Delta_{m,k+1}^{+}\rVert\bigr{]}$ (17)
$\leq Cr_{k}+\eta C\underbrace{\mathbb{E}\lVert Z_{m,k}\rVert}_{\leq\sqrt{\tau}\sigma_{0}\sqrt{\tfrac{M+1}{M}}}+\lambda\!\left(1+\underbrace{\mathbb{E}\lVert\bar{u}_{k+1}\rVert}_{\leq 1/\sqrt{M}}\right)$
$=Cr_{k}+\beta+\gamma,$ (18)
where
$\beta=\eta C\sqrt{\tau}\sigma_{0}\sqrt{\tfrac{M+1}{M}},\qquad\gamma=\lambda\!\left(1+\tfrac{1}{\sqrt{M}}\right)$
Iterating (18) for $k$ steps yields
$r_{k}$ $\leq C^{k}r_{0}+(\beta+\gamma)\sum_{j=0}^{k-1}C^{j}$
$=C^{k}r_{0}+\frac{\beta+\gamma}{1-C}\bigl{(}1-C^{k}\bigr{)}=C^{k}r_{0}+\frac{\beta+\gamma}{\alpha}\bigl{(}1-C^{k}\bigr{)}.$ (19)
Because $r_{0}=0$ and $C^{k}\!\to 0$, taking the limit in (19) with respect to $k$ gives
$\limsup_{k\to\infty}r_{k}\leq\frac{\beta+\gamma}{\alpha}=\frac{\eta(1-\alpha)\sqrt{\tau}\sigma_{0}\sqrt{\tfrac{M+1}{M}}}{\alpha}+\frac{\lambda}{\alpha}\!\left(1+\tfrac{1}{\sqrt{M}}\right).$ (20)
Consequently, with diminishing learning rate $\eta\to 0$ and many workers $M\gg 1$, we obtain
$\boxed{\lim_{k\to\infty}\mathbb{E}\bigl{[}\lVert\Delta_{m,k}^{+}\rVert\bigr{]}=\frac{\lambda}{\alpha}+\mathcal{O}\!\Bigl{(}\eta\sigma_{0}+\tfrac{1}{\sqrt{M}}\Bigr{)}}$
which completes the proof.
Tightness of the bound.
We write the exact recursion as $r_{k+1}=Cr_{k}+\lambda+\xi_{k}$ with $\mathbb{E}[\xi_{k}]=0$ and $|\xi_{k}|\leq\eta_{k}C\sqrt{\tau}\sigma_{0}\sqrt{(M+1)/M}+\lambda/\sqrt{M}$. Choosing a diminishing step size $\eta_{k}\!\to\!0$ or letting $M\!\to\!\infty$ makes $\xi_{k}\!\xrightarrow{k\to\infty}\!0$. Since $|r_{k+1}-\lambda/\alpha|\leq C^{k+1}|r_{0}-\lambda/\alpha|+\sum_{j=0}^{k}C^{j}|\xi_{k-j}|$, both terms on the right vanish and the sequence converges:
$\boxed{\displaystyle\lim_{k\to\infty}\mathbb{E}\bigl{[}\lVert\Delta_{m,k}^{+}\rVert\bigr{]}=\frac{\lambda}{\alpha}.}$
Thus the upper bound is tight. ∎
Theorem 4 (Monotone PAC-Bayes Gap Tightening with Valley Radius).
[…]
$\mathbb{E}_{x\sim Q_{r_{j}}}\![L_{\mathcal{D}}(x)]\leq\mathbb{E}_{x\sim Q_{r_{j}}}\![L_{S}(x)]+\underbrace{\sqrt{\frac{\,\tfrac{d}{2}(c_{r_{j}}-1-\log c_{r_{j}})+\frac{D_{0}}{2\sigma_{0}^{2}r_{j}^{1-\beta}}+\log\frac{nJ}{\delta}}{2(n-1)}}}_{\displaystyle\textstyle\text{gap}(r_{j})}.$
because $1-\beta>0$, $\text{gap}(r_{j+1})<\text{gap}(r_{j})$ for every consecutive pair in $\mathcal{G}$.
Proof.
We base our starting bound on the analysis carried out in chatterji2019intriguing and SAM . Following their footsteps, we use the PAC-Bayes bound derived for DNNs by dziugaite2017computing by building upon mcallester1999pac . Then, for any prior distribution $P$ over parameters in $d$ dimensional space, and for any posterior distribution $Q$, the following generalization guarantee holds with probability at least $1-\delta$ over the random draw of the training set $S$ with $n$ samples:
$\mathbb{E}_{x\sim Q}\![L_{\mathcal{D}}(x)]\leq\mathbb{E}_{x\sim Q}\![L_{S}(x)]+\sqrt{\frac{KL(Q||P)+\log\frac{n}{\delta}}{2(n-1)}}.$ (21)
If we assume that prior $P$ and posterior $Q$ are isotropic Gaussians, i.e. $P\sim\mathcal{N}(\mu_{P},\sigma_{P}^{2}I_{d})$ and $Q\sim\mathcal{N}(\mu_{Q},\sigma_{Q}^{2}I_{d})$ we can simply express the $KL$ divergence as follows:
$KL(Q||P)=\frac{d}{2}\left(\frac{\sigma_{Q}^{2}}{\sigma_{P}^{2}}-1+\log\frac{\sigma_{P}^{2}}{\sigma_{Q}^{2}}\right)+\frac{\|\mu_{Q}-\mu_{P}\|_{2}^{2}}{2\sigma_{P}^{2}}$
Let $r$ be the valley width which asymptotically converges to the ratio between push and pull force strengths as characterized by Theorem 3, i.e. $r=\frac{\lambda}{\alpha}$. Since, $\lambda$ and $\alpha$ hyperparameters are set without seeing the data, we can shape the prior variance based on $r$. Let us consider the following isotropic Gaussians for prior and posterior: $P\sim\mathcal{N}(0,r\sigma_{0}^{2}I_{d})$ and $Q\sim\mathcal{N}(\mu,cr\sigma_{0}^{2}I_{d})$ where $c$ is the data-dependent coefficient that impacts the posterior, we can re-write the PAC-Bayes bound in 21 as follows:
$\mathbb{E}_{x\sim Q}\![L_{\mathcal{D}}(x)]\leq\mathbb{E}_{x\sim Q}\![L_{S}(x)]+\sqrt{\frac{\frac{d}{2}\left(c-1-\log c\right)+\frac{\|\mu\|_{2}^{2}}{2r\sigma_{0}^{2}}+\log\frac{n}{\delta}}{2(n-1)}}.$
It may be tempting to conclude that increasing $r$ tightens the generalization gap by $1/\sqrt{r}$ based on the expression on the right-hand side above. However, our current assumptions do not account for how $\|\mu\|_{2}^{2}$ changes with increased prior variance. Without additional guarantees on the behavior of $\|\mu_{Q}\|_{2}^{2}$, such a claim would be misleading or incomplete.
To address this incompleteness, we employ the Langford-Caruana grid. We start by declaring a geometric grid of valley widths $\mathcal{G}$:
$\mathcal{G}=\bigl{\{}r_{j}=r_{\min}(1+\gamma)^{j}\big{|}j=0,\dots,J\bigr{\}},\qquad J=\Bigl{\lceil}\log_{1+\gamma}\!\frac{r_{\max}}{r_{\min}}\Bigr{\rceil},$
with $r_{\min},r_{\max},\gamma>0$ chosen a priori. For each $r_{j}$ we attach a Gaussian prior $P_{r_{j}}=\mathcal{N}(0,r_{j}\sigma_{0}^{2}I_{d})$. Now, running the algorithm with the target width $r_{j}$ yields the following posterior $Q_{r_{j}}=\mathcal{N}\bigl{(}\mu_{r_{j}},c_{r_{j}}r_{j}\sigma_{0}^{2}I_{d}\bigr{)}$ where $c_{r_{j}}\geq 1$ is again data-dependent.
Applying (21) to every $(P_{r_{j}},Q_{r_{j}})$ pair from the grid and union-bounding over the $J$ choices as practiced in Langford–Caruana method langford2001caruna gives, with probability at least $1-\delta$:
$\mathbb{E}_{x\sim Q_{r_{j}}}\bigl{[}L_{\mathcal{D}}(x)\bigr{]}\leq\mathbb{E}_{x\sim Q_{r_{j}}}\bigl{[}L_{S}(x)\bigr{]}+\underbrace{\sqrt{\frac{\tfrac{d}{2}\bigl{(}c_{r_{j}}-1-\log c_{r_{j}}\bigr{)}+\dfrac{\lVert\mu_{r_{j}}\rVert_{2}^{2}}{2r_{j}\sigma_{0}^{2}}+\log\tfrac{nJ}{\delta}}{2(n-1)}}}_{\text{gap}(r_{j})}$ (22)
We assume bounded-drift that sets an upper limit on the $L2$ norm of the posterior mean based on a chosen valley width. Assume that for some constants $D_{0}\geq 0$ and $0\leq\beta<1$
$\|\mu_{r}\|_{2}^{2}\leq D_{0}r^{\beta},\qquad\text{for all }r\in[r_{\min},r_{\max}].$ (23)
Substituting (23) into the numerator of the $\text{gap}(r_{j})$ term in (22) yields
$\text{gap}(r_{j})=\sqrt{\frac{\tfrac{d}{2}\bigl{(}c_{r_{j}}-1-\log c_{r_{j}}\bigr{)}+\dfrac{D_{0}}{2r_{j}^{1-\beta}\sigma_{0}^{2}}+\log\tfrac{nJ}{\delta}}{2(n-1)}}.$
Because $1-\beta>0$, the term $D_{0}/(r_{j}^{1-\beta})$ is strictly decreasing in $r_{j}$; all other terms are independent of $r_{j}$. Hence $\text{gap}(r_{j})$ is monotonically decreasing along the grid $\mathcal{G}$:
$r_{j+1}>r_{j}\Longrightarrow\text{gap}(r_{j+1})<\text{gap}(r_{j}).$
Additionally, since (22) holds simultaneously for every $r_{j}$, we can safely pick
$r^{\star}=\arg\min_{r_{j}\in\mathcal{G}}\text{gap}(r_{j})$
after training (or equivalently, choose the model with the smallest validation loss). The PAC-Bayes guarantee remains valid and satisfies $B(r^{\star})\leq B(r_{0})$.
E.3 Maximizing the Worker Consensus Distance in a Valley with Radius $C$
Consider $M$ points placed on a circle of radius $C$ in the 2D plane. Let each point $P_{i}$ be parameterized by an angle $\theta_{i}$, so that:
$P_{i}=(C\cos\theta_{i},C\sin\theta_{i}),\quad i=1,2,\ldots,M$
We want to maximize the mean distance of these points to the average variable, $P_{A}=(x_{A},y_{A})$, where the coordinates of the average variable can be expressed as follows:
$x_{A}=\frac{1}{M}\sum_{i=1}^{M}C\cos\theta_{i},\quad y_{A}=\frac{1}{M}\sum_{i=1}^{M}C\sin\theta_{i}$
The Euclidean distance $d_{i}$ from each point $P_{i}$ to the average $P_{a}$ can be written as:
$d_{i}=\sqrt{(C\cos\theta_{i}-x_{A})^{2}+(C\sin\theta_{i}-y_{A})^{2}}$
Recall that Simplified MV, is expressed as follows:
$R=\frac{1}{M}\sum_{i=1}^{M}d_{i}.$
Directly optimizing the above objective is harder due to the square root in $d_{i}$. Hence, we simplify the problem and maximize the following objective:
$R=\sum_{i=1}^{M}d_{i}^{2}.$
We plug in the expression for $d_{i}$ and re-write $R$:
$R=\sum_{i=1}^{M}\left((C\cos\theta_{i}-x_{A})^{2}+(C\sin\theta_{i}-y_{A})^{2}\right)$
If we expand each term in the summation, we obtain:
$R=\sum_{i=1}^{M}[C^{2}\cos^{2}\theta_{i}+C^{2}\sin^{2}\theta_{i}-2C\cos\theta_{i}x_{A}-2C\sin\theta_{i}y_{A}+x_{A}^{2}+y_{A}^{2}]$
Now, using the identity $\cos^{2}\theta_{i}+\sin^{2}\theta_{i}=1$, we can get:
$MC^{2}-2C\left(x_{A}\sum_{i=1}^{M}\cos\theta_{i}+y_{A}\sum_{i=1}^{M}\sin\theta_{i}\right)+M(x_{A}^{2}+y_{A}^{2})$
Since $x_{A}=\frac{C}{M}\sum_{i=1}^{M}\cos\theta_{i}$ and $y_{A}=\frac{C}{M}\sum_{i=1}^{M}\sin\theta_{i}$, we can rewrite the expression above as:
$MC^{2}-2C\left(x_{A}\cdot\frac{Mx_{A}}{C}+y_{A}\cdot\frac{My_{A}}{C}\right)+M(x_{A}^{2}+y_{A}^{2})$
Further simplification yields:
$S=MC^{2}-M(x_{A}^{2}+y_{A}^{2})$
This now tells us that maximizing $R$ is equivalent to minimizing $x_{A}^{2}+y_{A}^{2}$, which can be written as follows:
$x_{A}^{2}+y_{A}^{2}=\left(\frac{C}{M}\sum_{i=1}^{M}\cos\theta_{i}\right)^{2}+\left(\frac{C}{M}\sum_{i=1}^{M}\sin\theta_{i}\right)^{2}$
To minimize $x_{A}^{2}+y_{A}^{2}$, we require:
$\sum_{i=1}^{M}\cos\theta_{i}=0\quad\text{and}\quad\sum_{i=1}^{M}\sin\theta_{i}=0,$
Although this does not have a unique solution, one straightforward solution is the symmetric distribution of the points on the circle so that the above expressions can be made equal to 0. This can be mathematically expressed as follows:
$\theta_{i}=\theta_{0}+\frac{2\pi(i-1)}{M},\quad i=1,2,\ldots,M$
Here $\theta_{0}$ is any fixed angle used as a reference for the symmetric arrangement. Also observe that, in this configuration, the maximum mean distance is $C$.
E.4 Convergence in Non-Convex Setting
Formulation. Let us have $M$ workers collaboratively process non-exclusive data shards to find a DNN model that attains the smallest loss on the training data:
$\min\limits_{x\in\mathcal{R}^{d}}f(x)\stackrel{{\scriptstyle\Delta}}{{=}}\frac{1}{M}\sum_{i=1}^{M}f_{i}(x)$
We define each $f_{i}(x)$ as $f_{i}(x)\stackrel{{\scriptstyle\Delta}}{{=}}\mathbb{E}_{\xi_{i}\sim\mathcal{D}_{i}}\left[F_{i}(x;\xi_{i})\right]$ where $\mathcal{D}_{i}$ is the portion of the data seen by worker $i$. We assume that each worker can locally observe unbiased, independent stochastic gradients similar to yu2019parallel . $g_{i}^{t}=\nabla F_{i}(x_{i}^{t-1};\xi_{i}^{t})$ and $\mathbb{E}_{\xi_{i}^{t}\sim\mathcal{D}_{i}}\left[g_{i}^{t}|\xi^{t-1}\right]=\nabla f_{i}(x_{i}^{t-1})$ where $t$ denotes the iteration index. We assume the following technical conditions for the non-convex convergence rate analysis:
  • We assume each function $f_{i}(x)$ to be L-smooth:
    $\|\nabla f_{i}(a)-\nabla f_{i}(b)\|\leq L\|a-b\|$
  • Bounded variance:
    $\mathbb{E}_{\xi_{i}\sim\mathcal{D}_{i}}\left[\|\nabla F_{i}(x;\xi_{i})-\nabla f_{i}(x)\|^{2}\right]\leq\sigma^{2}$
  • Bounded domain at any iteration $t$:
    $\mathbb{E}\left[\|x_{i}^{t}-x_{A}^{t}\|^{2}\right]\leq\Delta^{2}$
    where $x_{A}^{t}$ is the average of the workers at iteration $t$, i.e. $x_{A}^{t}=\frac{1}{M}\sum_{i=1}^{M}x_{i}^{t}$.
Recall that the optimization objective we have is expressed as follows:
$\sum_{m=1}^{M}f_{i}(x_{i})+\frac{\alpha}{2}\|x_{i}-x_{C}\|^{2}-\frac{\lambda}{M}\sum_{i=1}^{M}\|x_{i}-x_{A}\|_{2}$
We consider a general update rule that corresponds to optimizing the objective above in a stochastic way.
$x_{i}^{t}=x_{t}^{t-1}-\eta g_{i}^{t}-\left(\alpha(x_{i}^{t-1}-x_{A}^{t-1})-\lambda\frac{x_{i}^{t-1}-x_{A}^{t-1}}{\|x_{i}^{t-1}-x_{A}^{t-1}\|}\right)$
Proof. We start by writing the following expression from $L$-smoothness assumption:
$\mathbb{E}\left[f(x_{i}^{t})\right]\leq\mathbb{E}$ $\left[f(x_{i}^{t-1})\right]$ (24)
$+$ $\mathbb{E}\left[\langle\nabla f(x_{i}^{t-1},x_{i}^{t}-x_{i}^{t-1})\rangle\right]$ (25)
$+$ $\frac{L}{2}\mathbb{E}\left[\|x_{i}^{t}-x_{i}^{t-1}\|^{2}\right]$ (26)
We individually tackle the terms on the right-hand side (RHS) of the inequality. We start with 26:
$\mathbb{E}$ $\left[\|x_{i}^{t}-x_{i}^{t-1}\|^{2}\right]$ (27)
$=\mathbb{E}$ $\left[\|-\eta g_{i}^{t}-\eta\alpha(x_{i}^{t-1}-x_{A}^{t-1})+\eta\lambda\frac{x_{i}^{t-1}-x_{A}^{t-1}}{\|x_{i}^{t-1}-x_{A}^{t-1}\|}\|^{2}\right]$
3$\leq$ 3≤ 3 $\mathbb{E}\left[\|-\eta g_{i}^{t}\|^{2}\right]$ (28)
$+$ $3\mathbb{E}\left[\|-\eta\alpha(x_{i}^{t-1}-x_{A}^{t-1})\|^{2}\right]$ (29)
$+$ $3\mathbb{E}\left[\|+\eta\lambda\frac{x_{i}^{t-1}-x_{A}^{t-1}}{\|x_{i}^{t-1}-x_{A}^{t-1}\|}\|^{2}\right]$ (30)
Where the inequality follows from $\|\sum_{i=1}^{n}a_{i}\|^{2}\leq n\sum_{i=1}^{n}\|a_{i}\|^{2}$ for $n=3$. We then individually consider the three terms on the RHS. We start by 28:
$\mathbb{E}\left[\|-\eta g_{i}^{t}\|^{2}\right]$
$=\eta^{2}\mathbb{E}\left[\|g_{i}^{t}-\nabla f_{i}(x_{i}^{t-1})+\nabla f_{i}(x_{i}^{t-1})\|^{2}\right]$
$\leq$ $\eta^{2}\left(\mathbb{E}\left[\|g_{i}^{t}-\nabla f_{i}(x_{i}^{t-1})\|^{2}\right]+\mathbb{E}\|\nabla f_{i}(x_{i}^{t-1})\|^{2}\right)$ (31)
$\leq$ $\eta^{2}\left(\sigma^{2}+\mathbb{E}\|\nabla f_{i}(x_{i}^{t-1})\|^{2}\right)$ (32)
Here the first inequality is obtained using $\mathbb{E}\left[\|A\|^{2}\right]=\mathbb{E}\left[\|A-\mathbb{E}\left[A\right]\|^{2}\right]+\|\mathbb{E}\left[A\right]\|^{2}$ and the second one is reached by using the bounded variance assumption. For 29 and 30, we have:
$\mathbb{E}$ $\left[\|-\eta\alpha(x_{i}^{t-1}-x_{A}^{t-1})\|^{2}\right]\leq\eta^{2}\alpha^{2}\Delta^{2}\hskip 8.53581pt\text{and}$
$\mathbb{E}$ $\left[\|+\eta\lambda\frac{x_{i}^{t-1}-x_{A}^{t-1}}{\|x_{i}^{t-1}-x_{A}^{t-1}\|}\|^{2}\right]\leq\eta^{2}\lambda^{2}$
By combining all these inequalities, we can derive an upper bound for 26 as follows:
$\frac{L}{2}\mathbb{E}$ $\left[\|x_{i}^{t}-x_{i}^{t-1}\|^{2}\right]$ (33)
$\leq\frac{3L}{2}$ $\eta^{2}\left(\sigma^{2}+\mathbb{E}\|\nabla f_{i}(x_{i}^{t-1})\|^{2}\right)$
$+$ $\frac{3L}{2}\eta^{2}\alpha^{2}\Delta^{2}+\frac{3L}{2}\eta^{2}\lambda^{2}$
We now derive an upper bound for 25.
$\mathbb{E}\langle\nabla f(x_{i}^{t-1},$ $x_{i}^{t}-x_{i}^{t-1})\rangle$ (34)
$=\mathbb{E}$ $[-\eta\nabla f(x_{i}^{t-1})^{T}(g_{i}^{t})]$ (35)
$+\mathbb{E}$ $[-\eta\alpha\nabla f(x_{i}^{t-1})^{T}(x_{i}^{t-1}-x_{A}^{t-1})$ (36)
$+\mathbb{E}$ $[\eta\lambda\nabla f(x_{i}^{t-1})^{T}\frac{x_{i}^{t-1}-x_{A}^{t-1}}{\|x_{i}^{t-1}-x_{A}^{t-1}\|}]$ (37)
For the first term on RHS, we can write the following by using $\mathbb{E}_{\xi_{i}^{t}\sim\mathcal{D}_{i}}\left[g_{i}^{t}|\xi^{t-1}\right]=\nabla f_{i}(x_{i}^{t-1})$.
$\mathbb{E}[-\eta\nabla f(x_{i}^{t-1})^{T}(g_{i}^{t})]=-\eta\mathbb{E}\|\nabla f_{i}(x_{i}^{t-1})\|^{2}$
Then using Young’s inequality, the upper bounds for the other two terms on the RHS can be written as follows:
$\mathbb{E}[-\eta\alpha\nabla$ $f(x_{i}^{t-1})^{T}(x_{i}^{t-1}-x_{A}^{t-1})]$
$\leq$ $\frac{\eta\alpha}{2}\mathbb{E}[\|\nabla f(x_{i}^{t-1})\|^{2}]+\frac{\eta\alpha}{2}\mathbb{E}[\|x_{i}^{t-1}-x_{A}^{t-1}\|^{2}]$
$\leq$ $\frac{\eta\alpha}{2}\mathbb{E}[\|\nabla f(x_{i}^{t-1})\|^{2}]+\frac{\eta\alpha}{2}\Delta^{2}$
Similarly for the second term, we have:
$\mathbb{E}[\eta\lambda\nabla$ $f(x_{i}^{t-1})^{T}\frac{x_{i}^{t-1}-x_{A}^{t-1}}{\|x_{i}^{t-1}-x_{A}^{t-1}\|})]$
$\leq$ $\frac{\eta\lambda}{2}\mathbb{E}[\|\nabla f(x_{i}^{t-1})\|^{2}]+\frac{\eta\lambda}{2}\mathbb{E}[||\frac{x_{i}^{t-1}-x_{A}^{t-1}}{\|x_{i}^{t-1}-x_{A}^{t-1}\|}||^{2}]$
$\leq$ $\frac{\eta\lambda}{2}\mathbb{E}[\|\nabla f(x_{i}^{t-1})\|^{2}]+\frac{\eta\lambda}{2}$
Overall, we can write the following for 25:
$\mathbb{E}\langle\nabla f(x_{i}^{t-1},$ $x_{i}^{t}-x_{i}^{t-1})\rangle$ (38)
$\leq$ $-\eta\mathbb{E}\|\nabla f_{i}(x_{i}^{t-1})\|^{2}$
$+\frac{\eta\alpha}{2}\mathbb{E}[\|\nabla f(x_{i}^{t-1})\|^{2}]+\frac{\eta\alpha}{2}\Delta^{2}$
$+\frac{\eta\lambda}{2}\mathbb{E}[\|\nabla f(x_{i}^{t-1})\|^{2}]+\frac{\eta\lambda}{2}$
Now using the upper bounds obtained in 38 and 33 we can write the following for 24:
$\mathbb{E}$ $[f(x_{i}^{t})]\leq\mathbb{E}[f(x_{i}^{t-1})]$
$-\eta\left(1-\frac{\alpha}{2}-\frac{\lambda}{2}-\frac{3L}{2}\eta\right)\mathbb{E}[\|\nabla f(x_{i}^{t-1})\|^{2}]$
$+\frac{3L}{2}\eta^{2}(\alpha^{2}\Delta^{2}+\lambda^{2}+\sigma^{2})+\frac{\eta}{2}(\alpha\Delta^{2}+\lambda)$
Furthermore, if we assume that $\left(1-\alpha-\lambda-3L\eta\right)>0$, we can also write:
$\frac{\eta}{2}\mathbb{E}$ $[\|\nabla f(x_{i}^{t-1})\|^{2}]\leq\mathbb{E}[f(x_{i}^{t-1})]-\mathbb{E}[f(x_{i}^{t})]$
$+\frac{3L}{2}\eta^{2}(\alpha^{2}\Delta^{2}+\lambda^{2}+\sigma^{2})+\frac{\eta}{2}(\alpha\Delta^{2}+\lambda)$
Finally, by multiplying both sides with $\frac{2}{\eta}$, summing the inequality from $t=0,1,...,T$ and taking the average of iterations and using $f(x^{*})\leq f(x^{T})$ where $x^{*}$ is the minimum of $f$, we reach the following:
$\frac{1}{T}\sum_{t=0}^{T}\mathbb{E}$ $[\|\nabla f(x_{i}^{t-1})\|^{2}]\leq\frac{2(f(x_{0})-f(x^{*}))}{\eta T}$
$+3L\eta(\alpha^{2}\Delta^{2}+\lambda^{2}+\sigma^{2})+\alpha\Delta^{2}+\lambda$
This characterizes the convergence rate of a single worker in a non-convex setting by bounding the expected value of gradient norms averaged over iterations.
Appendix F More Loss and Error Landscape Visualizations
Here we present more visualizations by including train loss, test loss, train error and test error landscapes obtained after training ResNet-18 models on CIFAR-10/100 datasets with SimpleAvg and DPPFSimpleAvg when the underlying local optimizer is SGD. The visualization settings—including whether the train or test set is used, whether loss or error is shown, the dataset, axis limits, and step sizes—are specified in the captions. Observe that in all the plots below, the optima recovered by DPPFSimpleAvg are substantially wider than those of vanilla SimpleAvg, and the corresponding test error and loss values are consistently lower.
We also present our visualization technique here, as we do not directly inherit a visualization method from the literature. To project the worker positions on a 2D plane with loss/error contours, we perform Singular Value Decomposition (SVD) among the distance vectors calculated between workers and the average variable. Then, we take the two most representative components (unit vectors) that capture the most variance. Using these vectors, we scan a 2D grid whose endpoints and resolution are specified, starting from the average variable $x_{A}$, hence the $x_{A}$ is always at the origin of the grid. We record the test and training loss and error statistics for each grid point, plot the resulting contours, and finally project the worker positions onto this plane. This procedure is described in Algorithm  3.
Input : trained model parameters from $M$ workers $x_{1},x_{2},...,x_{M}$; grid limit $L$ for defining the grid edges; step size $s$ for the resolution of the grid
$x_{A}\leftarrow\mathbf{0}$
for $m=1$ to $M$; do
 $x_{A}\leftarrow x_{A}+\frac{x_{m}}{M}$
end for
$\Delta\leftarrow[\hskip 5.69054pt]$
for $m=1$ to $M$; do
 $\Delta\leftarrow\Delta+[(x_{m}-x_{A})]$
end for
[…]
$\Phi\leftarrow[\hskip 5.69054pt]$
for $i=-L:s:L$; do
 for $j=-L:s:L$; do
    $x\leftarrow x_{A}+i\delta_{x}+j\delta_{y}$
    $\Phi\leftarrow\Phi+[f(x,D_{train}),f(x,D_{test})]$
   end for
end for
[…]
Algorithm 3 Landscape Visualization
(a) SimpleAvg
(b) DPPFSimpleAvg
[…]
(a) SimpleAvg
(b) DPPFSimpleAvg
[…]
(a) SimpleAvg
(b) DPPFSimpleAvg
[…]
(a) SimpleAvg
(b) DPPFSimpleAvg
[…]
(a) SimpleAvg
(b) DPPFSimpleAvg
[…]
(a) SimpleAvg
(b) DPPFSimpleAvg
[…]
(a) SimpleAvg
(b) DPPFSimpleAvg
[…]
(a) SimpleAvg
(b) DPPFSimpleAvg
[…]
(a) SimpleAvg
(b) DPPFSimpleAvg
[…]
(a) SimpleAvg
(b) DPPFSimpleAvg
[…]
(a) SimpleAvg
(b) DPPFSimpleAvg
[…]
(a) SimpleAvg
(b) DPPFSimpleAvg
[…]
(a) SimpleAvg
(b) DPPFSimpleAvg
[…]
(a) SimpleAvg
(b) DPPFSimpleAvg
[…]
(a) SimpleAvg
(b) DPPFSimpleAvg
[…]
(a) SimpleAvg
(b) DPPFSimpleAvg
[…]
(a) SimpleAvg
(b) DPPFSimpleAvg
[…]
(a) SimpleAvg
(b) DPPFSimpleAvg
[…]
(a) SimpleAvg
(b) DPPFSimpleAvg
[…]
(a) SimpleAvg
(b) DPPFSimpleAvg
[…]
(a) SimpleAvg
(b) DPPFSimpleAvg
[…]
(a) SimpleAvg
(b) DPPFSimpleAvg
[…]
(a) SimpleAvg
(b) DPPFSimpleAvg
[…]
(a) SimpleAvg
(b) DPPFSimpleAvg
[…]
(a) SimpleAvg
(b) DPPFSimpleAvg
[…]
(a) SimpleAvg
(b) DPPFSimpleAvg
[…]
(a) SimpleAvg
(b) DPPFSimpleAvg
[…]
(a) SimpleAvg
(b) DPPFSimpleAvg
[…]
(a) SimpleAvg
(b) DPPFSimpleAvg
[…]
(a) SimpleAvg
(b) DPPFSimpleAvg
[…]
(a) SimpleAvg
(b) DPPFSimpleAvg
[…]
(a) SimpleAvg
(b) DPPFSimpleAvg
[…]